### v3

# BotArtist: Generic approach for bot detection in Twitter via semi-automatic machine learning pipeline

Abstract
Twitter, as one of the most popular social networks, provides a platform for communication and online discourse. Unfortunately, it has also become a target for bots and fake accounts, resulting in the spread of false information and manipulation. In this paper, we introduce a semi-automatic machine learning pipeline (SAMLP) designed to address the challenge of detecting Twitter bots. Through this pipeline, we have developed a comprehensive bot detection model named BotArtist, based on user profile features. SAMLP leverages nine distinct publicly available datasets to train the BotArtist model. To assess BotArtist’s performance against current state-of-the-art solutions, we have selected 34 existing bot detection methods, each utilizing a diverse range of features. Our comparative evaluation of BotArtist and these existing methods, conducted across nine public datasets under standardized conditions, reveals that the proposed model outperforms existing solutions by almost 10% in terms of F1-score achieving an average score of 83.19 and 68.5 over specific and general approaches respectively. Following the acceptance of this paper, we will make both the developed SAMLP pipeline and the BotArtist model publicly accessible.
Introduction
Online social media has become an essential part of everyday life. During the past decade, online social platforms have managed to transform the communication routine of our daily lives. Due to their growing popularity, online social media gained millions of daily active users that not only consume the information but also create a space for content creators. The main reason behind online social media is real-time access to unlimited information, where registered users can share their comments and personal opinions about popular topics. Such high interest in online social phenomena generates the opportunity and the need for different categories of human interactions and content-sharing platforms. These opportunities lead to the creation of different online social platforms each of which provides a unique user experience, with a very similar goal of real-time human communication and content sharing. Twitter, one of the most popular social networks, with millions of active users, is used for news dissemination, political discussions, and social interactions. However, the platform has also been plagued by the presence of bots and fake accounts used to manipulate and spread false information. According to the research community, the usage of manipulation techniques implemented with the use of bot accounts is registered during diverse popular topic discussions. More specifically, studies show that bot accounts are involved in discussions around the 2016 and 2020 US Presidential elections (Golovchenko et al. 2020; Badawy, Ferrara, and Lerman 2018; Howard, Kollanyi, and Woolley 2016; Shevtsov et al. 2022, 2023b). Besides the US elections a high bot activity with spreading of misleading information is also detected during election periods (presidential/parliamentary/state) in countries like Germany, Sweden, France, Spain, etc. (Neudert, Kollanyi, and Howard 2017; Pastor-Galindo et al. 2020; Bradshaw et al. 2017; Fernquist, Kaati, and Schroeder 2018; Castillo et al. 2019; Rossi et al. 2020). Furthermore, election discourse is not the only point of interest for bot account creators. Spreading propaganda, advertisement, and fake news are identified during the vaccination debate (Broniatowski et al. 2018), the advertisement of e-cigarette (Allem and Ferrara 2016; Allem et al. 2017) and more recent examples of COVID-19 pandemic (Shahi, Dirkson, and Majchrzak 2021; Ferrara 2020; Yang et al. 2021). This high activity of bot accounts raises concerns in the research community and online social media platforms about the integrity of the shared information. Twitter has been taking measures to detect and suspend these bots to maintain the integrity of the platform. Most of the published studies utilize advanced ML techniques to analyze the content and features (structural/network).
Currently, the research community offers a variety of machine learning and neural network-based bot detection methods. Each of these provided methods excels at addressing specific bot detection scenarios, yielding optimal solutions for particular use cases. Unfortunately, the performance of these provided methods significantly diminishes in more general bot detection scenarios, encompassing different periods, discussion topics, and languages, among others. Recent studies have demonstrated that existing methods still fall short of achieving flawless Twitter bot detection, regardless of the number of user characteristics they incorporate.
In our research, we address this challenge by developing a semi-automatic machine learning pipeline (SAMLP) for constructing a generic bot detection model. Next, we compare this developed model with 34 currently available state-of-the-art approaches using nine publicly available datasets. This research yields two distinct evaluation scenarios: one specific to each dataset, where each model is trained and tested on each dataset separately, and a more generic approach where the models are trained and tested on the combined dataset comprising all nine unique datasets.
As a result of our research, we introduce BotArtist, a finely tuned general bot detection model that achieves the highest F1-score performance on three out of the nine datasets, with an average F1-score performance of 83.19. Additionally, the presented approach achieves greater accuracy in the general evaluation, with an overall F1-score improvement of more than 9% and an average improvement of 5% compared to the best-performing existing bot detection approaches. Moreover, BotArtist relies on only a limited set of profile features, allowing for the prediction of historical data independently of the Twitter API.
Finally, we offer insights into the decision-making process of BotArtist through the use of the SHAP method. This approach enables us to explain the model’s decisions and highlight the distinctions between normal and bot accounts as identified by the developed model. Upon paper acceptance, we will make the SAMLP pipeline and the generated BotArtist model publicly accessible.
Related Work
Bot detection on Twitter poses a formidable challenging task, primarily due to the increasing sophistication of these bots. Research efforts in this domain can generally be categorized into three main approaches: feature-based, text-based, and graph-based techniques. Each category offers a distinct angle on bot detection, leveraging various extracted characteristics of users and their activities on the social media platform.
Feature based
Methods falling into this category perform feature engineering based on information extracted from Twitter user profiles and their activity patterns. Researchers utilize traditional machine learning or neural network classification algorithms for bot detection. These approaches employ different sets of user characteristics as feature sets, including user metadata (Kudugunta and Ferrara 2018), tweets (Miller et al. 2014), description (Hayawi et al. 2022), temporal patterns (Mazza et al. 2019), and follow relationships (Feng et al. 2021a). Moreover, some aim to enhance the scalability of feature-based methods (Yang et al. 2020), discover unknown bot accounts using correlation-based techniques (Chavoshi, Hamooni, and Mueen 2016), and improve the trade-off between precision and recall in bot detection (Morstatter et al. 2016). However, creators of bot accounts are increasingly aware of the features utilized by the research community, leading to novel bot implementations designed to evade detection based on known features (Cresci 2020). Consequently, existing feature-based methods face several challenges in the accurate detection of these new bot accounts (Feng et al. 2021a).
Text based
Text-based methods rely on natural language processing (NLP) techniques for Twitter bot detection, extracting characteristics from posted content (tweets) and user profile descriptions. Approaches in this category include word embeddings (Wei and Nguyen 2019), recurrent neural networks (RNN) (Kudugunta and Ferrara 2018), attention mechanisms (Feng et al. 2021a), and the adoption of pre-trained language models for encoding tweets (Dukić, Keča, and Stipić 2020). Researchers have also combined tweet representations with user profile characteristics (Cai, Li, and Zengi 2017), employed unsupervised machine learning models (Feng et al. 2021a), and addressed multilingual context issues (Knauth 2019). Despite the extensive existing research studies and impressive performances in text-based approaches, novel bot accounts can still evade detection by sharing stolen content from genuine users (Cresci 2020). Additionally, recent work has demonstrated that relying solely on textual information is insufficient for robust and accurate bot detection (Feng et al. 2021c).
Graph based
The graph-based category of bot detection methods combines geometric deep neural networks with graph analytics. Current implementations leverage techniques such as node centrality (Dehghan et al. 2023), node representation learning (Pham et al. 2022), graph neural networks (GNNs) (Ali Alhosseini et al. 2019), and heterogeneous GNNs (Feng et al. 2021c) to conduct graph-based bot detection. In recent research, authors have borrowed ideas from other categories to merge different approaches, such as combining graph and text-based methods (Guo et al. 2021b; Yang, Ferrara, and Menczer 2022). Furthermore, novel GNN architectures have been proposed to exploit the heterogeneity in the Twitter network (Feng et al. 2022a). Generally, these approaches hold significant promise for Twitter bot detection.
Unfortunately, existing Twitter bot detection methods have their limitations. Simplified feature-based approaches struggle to generalize and provide robust results, while more complex methods based on text or graph characteristics require intensive computational resources and large datasets for model development. These limitations, coupled with the recent announcement of the monetizing Twitter API (Twitter 2023), consist of many existing methods expensive to maintain or nearly impossible to operate daily. In this research, we address these challenges and present robust and accurate Twitter bot detection methods that rely on a lightweight set of features. Our methods require only a single user profile object request and can predict users using historical user objects (Twitter API v1.1 or v2) without additional Twitter API requests, thereby significantly reducing operational costs.
Datasets
For this research paper, we have collected nine well-known publicly available datasets (Feng et al. 2021b). All the selected datasets already contain ground truth labels, primarily obtained through manual analysis or crowd-sourcing. To simplify, we have labeled the selected datasets as follows: C-15 (Cresci et al. 2015), G-17 (Gilani et al. 2017), C-17 (Cresci et al. 2017b, a), M-18 (Yang et al. 2020), C-S-18 (Cresci et al. 2018, 2019), C-R-19 (Mazza et al. 2019), B-F-19 (Yang et al. 2019), TwiBot-20 (Feng et al. 2021b), and TwiBot-22 (Feng et al. 2022b). In table: 1, we present the information provided in each dataset, along with the volume of normal and bot accounts.
Dataset C-15 G-17 C-17 M-18 C-S-18 C-R-19 B-F-19 TwiBot-20 TwiBot-22
# Total User 5,301 2,484 14,368 50,538 13,276 693 518 229,580 1,000,000
# Human 1,950 1,394 3,474 8,092 6,174 340 380 5,237 860,057
# Bot 3,351 1,090 10,894 42,446 7,102 353 138 6,589 139,943
# Total Tweet 2,827,757 0 6,637,615 0 0 0 0 33,488,192 88,217,457
# Human Tweet 2,631,730 0 2,839,361 0 0 0 0 33,488,192 81,250,102
# Bot Tweet 196,027 0 3,798,254 0 0 0 0 33,488,192 6,967,355
# Graph Edges 7,086,134 0 6,637,615 0 0 0 0 33,488,192 170,185,937
Table 1: Description of the selected datasets and the information contained in those datasets.
Figure 1: Machine learning pipeline used for BotArtist model creation. Includes feature selection and model fine-tuning; each step is executed during separate K-Fold cross-validation.
Methodology
The proposed approach is based on the development of a semi-automatic machine learning pipeline (SAMLP). We have chosen this implementation due to its simplicity for further usage and its ability to avoid many trivial mistakes during data processing, feature selection, hyper-parameter fine-tuning, and model evaluation. The steps of the developed pipeline are presented in figure: 1, where the initial step involves data splitting.
During data splitting, it is crucial to maintain a class balance between the training/validation and testing data portions. For this purpose, we utilize a stratified data split with a ratio of 70% for training/validation and 30% for the testing data portions, respectively. This data separation allows us to keep exactly 70% of each class during the training, while preserving the natural class distribution differences, present in the dataset. The testing data portion is kept hidden and is only utilized during the final testing of the model.
Feature Selection
After the data splitting, our pipeline applies a feature selection procedure over the train/validation data portion. For this purpose, we use the K-Fold cross-validation approach, which allows us to utilize the same data portion for multiple trainings and evaluations. Although K-Fold cross-validation may lead to slightly over-optimistic performance estimations, in our case, it is not relevant since we use this procedure for alpha hyper-parameter fine-tuning of the Lasso model, and we are interested in identifying the best alpha parameter.
Additionally, since the Lasso regression model does not perform well over imbalanced datasets, we also take into account the class imbalance of the train/validation dataset. When the class balance is not equal, we utilize under-sampling of the majority class, making the data perfectly balanced and improving the prediction performance of the Lasso model. Due to the high-class imbalance and the under-sampling of the majority class, there is a very high probability of losing some important samples, which may lead to inaccurate feature selection. To address this, we utilize 10 executive repetitions to capture as many samples of the majority class as possible.
During these repetition rounds, we store the best alpha parameters based on the square mean error metric from the entire stratified k-Fold cross-validation results, where $K=5$. At the end of the procedure, we compute the most frequent best $\alpha$ parameter and train the Lasso model with the entire train/validation dataset, keeping features with non-zero coefficients. Additionally, we check if the selected $\alpha$ parameter falls within the defined limits of the searching area. If it falls on the minimum or maximum values, we create a new searching area based on the original site to ensure the selected $\alpha$ is genuinely the best.
This approach allows us to utilize feature selection without additional knowledge of the original data and modifications, as the procedure automatically finds the best alpha values, manages class balance and imbalance via under-sampling and repeated executions, and processes both perfectly balanced and highly imbalanced datasets without information loss.
Model Fine-Tuning
Having reduced the problem’s dimensionality through feature selection, we now focus on identifying the model and its configuration (hyper-parameter set) that can provide accurate predictions over unseen data samples. We selected three well-known machine learning classification models: SVM (Boser, Guyon, and Vapnik 1992), RandomForest (Breiman 2001), and state-of-the-art XGBoost (Extreme Gradient Boosting) (Chen et al. 2015).
For each model, we defined large hyper-parameter ranges to cover a variety of configurations. To reduce complexity and execution time, we developed a sampling method that randomly selects $C$ unique configurations per model (in our case, $C=50$). This method allows us to estimate different possible configurations without sacrificing execution time. Additionally, since the dataset may be affected by class imbalance, we compute class-specific weights to construct selected classifiers with consideration of class imbalance.
To evaluate each selected configuration, we use a stratified K-Fold cross-validation approach, allowing us to utilize the entire train/validation data portion and evaluate each configuration over multiple validation data portions. During the evaluation, we compute the average configuration performance measured in the F1-score, which can effectively measure model performance over both binary and multi-class classification tasks, considering class imbalance.
Afterward, we select the best classifier and configuration based on the average validation (F1-score). This model configuration becomes the final model, which will be evaluated over the hold-out (testing) data portion. In the final stage of the model development, we provide model explainability using the SHAP game-theoretical approach (Shapley 1953). This allows us to describe the reasons behind the model’s decisions and highlight the distinctions between classes.
Additionally, in the case of binary classification, we need to adjust the decision threshold, as binary classification models are not perfectly aligned with the 50% decision threshold. To do this, we utilize the precision vs. recall curve to identify the decision threshold that maximizes the model’s performance. The developed model is then trained over the entire dataset, including the train/validation and testing sets, and can be utilized for real-world applications with the identified decision threshold.
Feature Type Feature Type calculation
name_len count screen_name_sim real-valued Jaccard of (screen_name, user_name)
screen_name_len count foll_friends real-valued follower / friends
description_len count age real-valued created_at / collection date
listed count listed_by_age real-valued listed / account age
statuses count statuses_by_age real-valued statuses / account age
followers count followers_by_age real-valued followers / account age
following count following_by_age real-valued friends / account age
name_upper_len count name_upper_pcnt real-valued percentage of upper case
name_lower_len count name_lower_pcnt real-valued percentage of lower case
name_digits_len count name_digits_pcnt real-valued percentage of digits
name_special_len count name_special_pcnt real-valued percentage of other characters
screen_name_upper_len count screen_name_upper_pcnt real-valued percentage of upper case
screen_name_lower_len count screen_name_lower_pcnt real-valued percentage of lower case
screen_name_digits_len count screen_name_digits_pcnt real-valued percentage of digits
screen_name_special_len count screen_name_special_pcnt real-valued percentage of other characters
description_upper_len count description_upper_pcnt real-valued percentage of upper case
description_lower_len count description_lower_pcnt real-valued percentage of lower case
description_digits_len count description_digits_pcnt real-valued percentage of digits
description_special_len count description_special_pcnt real-valued percentage of other characters
description_urls count name_entropy real-valued entropy of user name
description_mentions count screen_name_entropy real-valued entropy of the screen name
description_hashtags count has_location boolean
total_urls count has_profile_image boolean
protected boolean has_profile_url boolean
verified boolean
Table 2: List of the profile features extracted from the Twitter API user objects.
Method Type C-15 G-17 C-17 M-18 C-S-18 C-R-19 B-F-19 TB-20 TB-22 Average
SGBot F 77.9 72.1 94.6 99.5 82.3 82.7 49.6 84.9 36.6 75.57
Kudugunta et al. F 75.3 49.8 91.7 94.5 50.9 49.2 49.6 47.3 51.7 62.22
Hayawi et al. F 85.6 34.7 93.8 91.5 60.8 60.9 20.5 77.1 24.7 61.06
BotHunter F 97.2 69.2 91.6 99.6 82.2 82.9 49.6 79.1 23.5 74.98
NameBot F 83.4 44.8 85.7 91.6 61.1 67.5 38.5 65.1 0.5 59.80
Abreu et al. F 76.4 66.7 95.0 97.9 76.9 83.5 53.8 77.1 53.4 75.63
BotArtist F 98.3 76.1 97.0 99.7 80.6 88.3 68.4 82.2 58.2 83.19
Cresci T 1.17 - 22.8 - - - - 13.7 - -
Wei T 82.7 - 78.4 - - - - 57.3 53.6 -
BGSRD T 90.8 35.7 86.3 90.5 58.2 41.1 13.0 70.0 21.1 56.30
RoBERTa T 95.8 - 94.3 - - - - 73.1 20.5 -
T5 T 89.3 - 92.3 - - - - 70.5 20.2 -
Efthimion FT 94.1 5.2 91.8 95.9 68.2 71.7 0.0 67.2 27.5 57.95
Kantepe FT 78.2 - 79.4 - - - - 62.2 58.7 -
Miller FT 83.8 59.9 86.8 91.1 56.8 43.6 0.0 74.8 45.3 60.23
Varol FT 94.7 - - - - - - 81.1 27.5 -
Kouvela FT 98.2 66.6 99.1 98.2 80.4 81.1 28.1 86.5 30.0 74.24
Santos FT 78.8 14.5 83.0 92.4 65.2 75.7 21.0 60.3 - -
Lee FT 98.6 67.8 99.3 97.9 82.5 82.7 50.3 80.0 30.4 76.61
LOBO FT 98.8 - 97.7 - - - - 80.8 38.6 -
Moghaddam FG 73.9 - - - - - - 79.9 32.1 -
Alhosseini FG 92.2 - - - - - - 72.0 38.1 -
Knauth FTG 91.2 39.1 93.4 91.3 94.0 54.2 41.3 85.2 37.1 69.64
FriendBot FTG 97.6 - 87.4 - - - - 80.0 - -
SATAR FTG 95.0 - - - - - - 86.1 - -
Botometer FTG 66.9 77.4 96.1 46.0 79.6 79.0 30.8 53.1 42.8 63.5
Rodrifuez-Ruiz FTG 87.7 - 85.7 - - - - 63.1 56.6 -
GraphHist FTG 84.5 - - - - - - 67.6 - -
EvolveBot FTG 90.1 - - - - - - 69.7 14.1 -
Dehghan FTG 88.3 - - - - - - 76.2 - -
GCN FTG 97.2 - - - - - - 80.8 54.9 -
GAT FTG 97.6 - - - - - - 85.2 55.8 -
HGT FTG 96.9 - - - - - - 88.2 39.6 -
SimpleHGN FTG 97.5 - - - - - - 88.2 45.4 -
BotRGCN FTG 97.3 - - - - - - 87.3 57.5 -
RGT FTG 97.8 - - - - - - 88.0 42.9 -
Table 3: The performance of each selected bot detection model, as reported in the (Feng et al. 2021b) paper, is compared with that of BotArtist. Performance is measured using the F1-score. In this benchmark, each model is trained and tested on each dataset separately.
Method C-15 G-17 C-17 M-18 C-S-18 C-R-19 B-F-19 TB-20 TB-22 Total Average
BotArtist 82.7 39.9 87.3 99.0 80.6 73.8 16.6 80.3 56.9 63.7 68.5
Lee 82.3 0.0 83.6 97.7 78.2 67.7 20.0 8.5 42.4 52.9 53.3
Abreu 84.4 0.3 80.1 88.4 67.1 40.8 11.7 15.6 29.0 40.4 46.3
SGBot 75.0 3.6 79.8 99.2 76.7 68.9 0.0 15.2 43.3 53.8 51.5
BotHunter 73.4 7.1 76.0 99.2 76.0 44.8 11.1 14.7 28.0 43.1 47.8
Kouvela 95.5 20.4 94.7 98.1 78.4 71.4 21.0 28.5 36.0 52.0 60.5
Botometer 66.9 77.4 96.1 46.0 79.6 79.0 30.8 53.1 42.8 45.3 63.5
Table 4: The measurement of usage, like general bot detection approaches, involves training and testing models on all datasets. Performance is assessed using the F1-score.
Feature Extraction
As mentioned earlier, in this study we pursue the creation of a simple model that in comparison with other bot detection systems does not require a large volume of information or features for accurate prediction. In our case, the number of Twitter profiles triggers our attention. The majority of social media bot accounts are created for content promotion and increasing the attention of other registered users. To maximize this goal, bot accounts pursue the expansion of social audience. Social expansion is very difficult to achieve without high activity, such as tweets, re-tweets, comments, etc. In comparison with profile-based features, the extraction and utilization of textual or graph features can result in instability, causing a significant performance drop across various periods or discussion topic datasets (Shevtsov et al. 2023a). Consequently, we need to extract as much information as possible from the user profiles to identify extremely active accounts. The difference between very popular social media accounts (also known as celebrities), and bot accounts is that they try to post and share as much content as possible during the day. To properly separate normal and bot accounts, we develop 49 unique features described in table: 2. Due to the differences between Twitter API v1.1 and v2, we have devised a method for constructing features in a way that allows the extraction of selected information from both v1.1 and v2 without any additional manipulation. The profile features we have extracted are categorized into three distinct groups: count, boolean, and real-valued.
The initial count feature categories encompass raw Twitter API object values, including followers, followings, status, and number of subscribed Twitter lists. In addition to these raw values, we have calculated statistics for textual fields such as user name, screen_name, and description. For these features, we have quantified the number of distinct characters, including uppercase letters, lowercase letters, digits, and special characters that do not fall into any of the previous categories. Furthermore, we have leveraged the profile description field, where users can provide additional information such as mentions of other users/organizations, hashtags, and additional URLs. In this context, we have counted the volume of the mentioned entities in the description field as raw values.
In contrast to the count categories, where values are represented as integers, the real-valued category contains floating-point values. We have conducted various comparisons and measurements across the profile fields in this category. Initially, we compute the age of the accounts, which allows us to distinguish highly active users with millions of posts spanning several years from users with similar activity but with an active period of only a few days. Account age is measured in days, calculated as the duration between the account creation date, provided by the Twitter API, and the collection date. Based on the account’s age, we can compute metrics similar to those in the count values category but adjusted for the account’s age perspective. For example, we compute metrics such as statuses_by_age, follower_by_age, and following_by_age using the formula:
$feature\_by\_age=\frac{feature\ value}{profile\ age}$
Given that many bot accounts are mass-created by automated scripts, they often have very similar or even trivial user names and screen names. Therefore, we have also applied the Jaccard similarity measure between user names and screen names. Jaccard similarity calculates the intersection of characters used in both fields in relation to the differences between them, using the following formula:
$Jaccard\ similarity(A,B)=\frac{|A\cap B|}{|A\cup B|}$
Additionally, we have computed the entropy of the user name and screen_name, enabling us to assess the randomness of these selected fields. In addition to the count values derived from the profile text fields (user name, screen_name, and description), we have computed the percentage of each specific character category, including lowercase, uppercase, digits, and special characters, relative to the length of the text field. These measurements offer a more precise means of identifying textual preferences more flexible to varying text lengths since they are computed as percentages of the field length.
The final category of extracted features consists of Boolean values, used to identify information that may or may not be provided by the account creator. For instance, we determine if the user account is protected, and verified, as well as extract information about the provided user location, profile URL, and whether the profile uses the default profile image.
By employing this comprehensive feature engineering approach, we have maximized feature extraction while working with the limited profile information available through Twitter API v2. In total, we have extracted 49 unique profile features, representing the full set of features used before the feature selection process in the BotArtist model.
Experimental Results
According to the methodology of SAMLP described earlier, we’ve successfully developed a semi-automated machine learning pipeline capable of constructing classification models (both binary and multi-class) for both balanced and imbalanced datasets. Using this pipeline, we created and tested our custom implementation called BotArtist for Twitter bot detection, utilizing nine well-known datasets. To evaluate BotArtist’s performance, in comparison to 34 other existing research approaches, we employed the most comprehensive Twitter bot detection benchmark available (Feng et al. 2022b), where various methods are compared using identical training/validation and testing data portions. This bench-marking allowed us to make a fair comparison between our approach and established bot detection methods.
In our comparison, we not only evaluate BotArtist against bot detection approaches based on similar categories of features but also compare it to existing approaches developed based on different sets of characteristics. The selected methods are categorized into five different groups.
Profile-Based features (F)
This category includes approaches that rely on features extracted from user profile objects, including SGBot (Yang et al. 2020), Kudugunta (Kudugunta and Ferrara 2018), Hayawi (Hayawi et al. 2022), BotHunter (Beskow and Carley 2018), NameBot (Beskow and Carley 2019) and Abreu (Abreu, Ralha, and Gondim 2020) implementations.
Text-Based features (T)
In this category, we have approaches that focus on text characteristics, including Cresci (Cresci et al. 2016), Wei (Wei and Nguyen 2019), BGSRD (Guo et al. 2021a), RoBERTa (Liu et al. 2019) and T5 (Raffel et al. 2020).
Profile and Text-Based features (FT)
These approaches combine both profile and text-based features, and the selected implementations are Efthimion (Efthimion, Payne, and Proferes 2018), Kantepe (Kantepe and Ganiz 2017), Miller (Miller et al. 2014), Varol (Varol et al. 2017), Kouvela (Kouvela, Dimitriadis, and Vakali 2020), Lee (Lee, Eoff, and Caverlee 2011) and LOBO (Echeverria et al. 2018).
Profile and Graph-Based features (FG)
Some approaches combine profile features with graph-based information. This category includes Moghaddam (Moghaddam and Abbaspour 2022) and Alhosseini (Ali Alhosseini et al. 2019) implementations.
Profile, Text, and Graph-Based features (FTG)
Approaches in this category aim to extract and combine as many features as possible, encompassing profile, text, and graph-based information. The selected implementations are Knauth (Knauth 2019), FriendBot (Beskow and Carley 2020), SATAR (Feng et al. 2021a), Botometer (Yang, Ferrara, and Menczer 2022), Rodríguez-Ruiz (Rodríguez-Ruiz et al. 2020), GraphHist (Magelinski, Beskow, and Carley 2020), EvolveBot (Yang, Harkreader, and Gu 2013), Dehghan (Dehghan et al. 2023), GCN (Kipf and Welling 2016), GAT (Veličković et al. 2017), HGT (Hu et al. 2020), SimpleHGN (Lv et al. 2021), BotRGCN (Feng et al. 2021c) and RGT (Feng et al. 2022a).
Models comparison
Our goal with BotArtist was to create a model with the highest possible level of generalizability. To achieve this, we designed two separate comparison scenarios. Initially, our focus was on assessing the fine-tuned BotArtist’s ability to capture general patterns within each dataset in our collection. As a result, we conducted training and testing for each of the selected models, including BotArtist, on each of the nine chosen datasets individually. This comparative approach allowed us to evaluate how well these selected models performed on each specific dataset, serving as a more rigorous bot detection implementation.
Additionally, the designed experiments provided us with a broader understanding of the overall capabilities of the selected methods when applied in more general scenarios. While implementing these experiments, we took into account that the performance would likely be higher than that of a generic approach. Therefore, our primary focus was on discerning the differences between the models to select the most accurate bot detection methods for the broader, more generalized comparison.
Previously, we had evaluated and compared models on each isolated dataset and calculated their average performance. While this approach provided us with an initial insight into model performance across diverse datasets, it lacked an assessment of model generalizability. To address this, we designed additional comparisons where we assessed the model’s ability to capture general patterns across different datasets. In these cases, we combined the training data from the nine datasets into a single dataset for model training. Combining data from multiple datasets, each with varying periods, discussion topics, and communities, demanded a high level of model generalization to avoid overfitting.
To measure the performance of the trained methods, we evaluated them by testing on portions of each dataset and also on the union of all the dataset testing portions. These approaches allowed us to not only identify the general performance of each bot detection method but also gain insights into performance variations across each specific dataset.
Parameter Value
max_depth
learning_rate
subsample
colsample_bytree
min_child_weight
gamma
reg_lambda
n_estimators
eval_metric
tree_method
predictor
objective
scale_pos_weight
selected features
Table 5: The following are the selected parameters of the final general BotArtist (XGBoost) model, along with the number of features chosen during the feature selection procedure.
In the initial model comparison experiment, we compared selected methods and assessed their performance based on the F1-score. This choice of metric was motivated by the highly imbalanced nature of the datasets, where proper comparison necessitates measurement metrics that account for class imbalances.
In table: 3, we present the results of this comparison. Notably, the BotArtist model emerged as the only bot detection method that achieved superior performance across three different datasets (M-18, C-R-19, B-F-19). Furthermore, the BotArtist model demonstrated the highest average F1-score, reaching 83.19. This represents a substantial improvement of 6.5% over the most accurate of the existing methods, showcasing the model’s capacity to accurately capture patterns within each distinct dataset.
As shown in table: 3, some datasets posed challenges for more complex methods that rely on text and graph features. These methods struggled to utilize datasets without text or graph information due to their limitations in accessing essential information required for the model.
In the second model comparison scenario, our objective was to assess the generalization capabilities of various models. To achieve this, we selectively choose the most accurate models from table: 3, in addition to the Botometer model (Yang, Ferrara, and Menczer 2022). The Botometer model was included because it had already been trained on a diverse range of datasets and was expected to perform well in general-case scenarios. Additionally, the SAMLP pipeline manages to fine-tune BotArtists with the parameters presented in the table: 5 where the XGBoost classification model achieves higher performance during the K-Fold cross-validation.
The results, as presented in table: 4, indicate that the most generalizable models are BotArtist, Botometer, and SGBot. BotArtist demonstrated superior performance over existing methods in terms of both F1-total and average scores, surpassing almost 10% existing methods in total. These results confirm that fine-tuned models which are based on the limited set of features are capable of capturing the general differences between normal and bot accounts.
Furthermore, it’s noteworthy that the BotArtist model outperformed existing methods in both comparison scenarios. Due to the high volume of experiments, we do not provide any information according to the selected features and final model configuration for each evaluation scenario. These data will be shared and described in detail, in a GitHub repository after the paper acceptance.
Figure 2: BotArtist model decision explainability figure produced by SHAP methodology.
Explainability
Besides the creation of the high-performance model used for bot detection, we aim to shed light on the model decision itself. For this purpose we employ SHapley Additive exPlanations (SHAP) values (Lundberg and Lee 2017) which offer various benefits. One of the most significant advantages of SHAP values is that they are not tied to any particular type of machine learning model and are therefore model-agnostic. Furthermore, SHAP values display local accuracy, and consistency, which are characteristics that are not found concurrently in other methods.
In this section we explain the BotArtist model trained over multiple datasets. Figure: 2 presents the 20 most important features and their impact on the final decision. The features are sorted in descending order of importance. Based on this order, we observe that the age of the account and the level of social activity are the most critical characteristics of our model, since the creation date of the account.
The presented figure showcases the values of the features, ranging from low (indicated by the blue color), to high (indicated by the red color). Simultaneously, the x-axis provides information on the impact of each feature value on the model decision. Based on the presented results, it is crucial to highlight that the SHAP explanation technique provides insights that align with past related works. For instance, we find that bot accounts tend to have a significantly lower age of the account and lower number of followers combined with a higher number of followers by age of the account. Based on this information we can assume that bot accounts are not so popular in terms of total followers in comparison with normal users, but the growth of the profile followers is higher in comparison with normal users. This assumption agrees with the original target of the bot account creators to achieve a higher audience during the short period.
Conclusions and Future Work
In this research paper, we introduce a semi-automatic machine learning pipeline (SAMLP) that addresses multiple challenges in creating machine learning models, including feature selection, hyperparameter fine-tuning, model evaluation, decision threshold optimization for binary classification, and provides model explainability through the SHAP game-theoretical approach. We apply this developed pipeline to create BotArtist, a versatile bot detection model based on user profile features. Our approach is trained and evaluated alongside current state-of-the-art solutions using nine different datasets. As demonstrated in our experiments, BotArtist surpasses existing, more complex methods in terms of generalization, achieving almost a 10% increase in total F1-score in multiple dataset comparisons and a 6.5% increase in separated dataset comparisons. Additionally, we offer insights into the final model’s decision-making process and the patterns it captures, based on the SHAP game-theoretical approach. Upon acceptance of this paper, we will make the developed SAMLP pipeline available for further research, along with the BotArtist model, which can label existing profile data. One of the limitations of the presented methodology is the relatively limited feature set, which might make it susceptible to evasion by future bot accounts that specifically aim to avoid detection based on these features. Further research conducted over a longer period is required to assess the model’s effectiveness in detecting bot accounts in the evolving landscape of social media.

### v4

@@
 Unfortunately, it has also become a target for bots and fake accounts, resulting in the spread of false information and manipulation.
-In this paper, we introduce a semi-automatic machine learning pipeline (SAMLP) designed to address the challenge of detecting Twitter bots.
-Through this pipeline, we have developed a comprehensive bot detection model named BotArtist, based on user profile features.
+This paper introduces a semi-automatic machine learning pipeline (SAMLP) designed to address the challenges correlated with machine learning model development.
+Through this pipeline, we develop a comprehensive bot detection model named BotArtist, based on user profile features.
 SAMLP leverages nine distinct publicly available datasets to train the BotArtist model.
-To assess BotArtist’s performance against current state-of-the-art solutions, we have selected 34 existing bot detection methods, each utilizing a diverse range of features.
-Our comparative evaluation of BotArtist and these existing methods, conducted across nine public datasets under standardized conditions, reveals that the proposed model outperforms existing solutions by almost 10% in terms of F1-score achieving an average score of 83.19 and 68.5 over specific and general approaches respectively.
-Following the acceptance of this paper, we will make both the developed SAMLP pipeline and the BotArtist model publicly accessible.
-Introduction
+To assess BotArtist’s performance against current state-of-the-art solutions, we select 35 existing Twitter bot detection methods, each utilizing a diverse range of features.
+Our comparative evaluation of BotArtist and these existing methods, conducted across nine public datasets under standardized conditions, reveals that the proposed model outperforms existing solutions by almost 10%, in terms of F1-score, achieving an average score of 83.19 and 68.5 over specific and general approaches respectively.
+As a result of this research, we provide a dataset of the extracted features combined with BotArtist predictions over the 10.929.533 Twitter user profiles, collected via Twitter API during the 2022 Russo-Ukrainian War, over a 16-month period.
+This dataset was created in collaboration with (Shevtsov et al., 2022a) where the original authors share anonymized tweets on the discussion of the Russo-Ukrainian war with a total amount of 127.275.386 tweets.
+The combination of the existing text dataset and the provided labeled bot and human profiles will allow for the future development of a more advanced bot detection large language model in the post-Twitter API era.
+1 Introduction
 Online social media has become an essential part of everyday life.
-During the past decade, online social platforms have managed to transform the communication routine of our daily lives.
-Due to their growing popularity, online social media gained millions of daily active users that not only consume the information but also create a space for content creators.
-The main reason behind online social media is real-time access to unlimited information, where registered users can share their comments and personal opinions about popular topics.
-Such high interest in online social phenomena generates the opportunity and the need for different categories of human interactions and content-sharing platforms.
-These opportunities lead to the creation of different online social platforms each of which provides a unique user experience, with a very similar goal of real-time human communication and content sharing.
+During the past decade, social networks have transformed the communication routine of our daily lives.
+Due to their growing popularity, online social media gained millions of daily active users not only consuming the information but also creating a space for content creators.
+The main reason behind the success of online social networks is the real-time access to unlimited information, where registered users can share their comments and personal opinions on popular topics.
 Twitter, one of the most popular social networks, with millions of active users, is used for news dissemination, political discussions, and social interactions.
-However, the platform has also been plagued by the presence of bots and fake accounts used to manipulate and spread false information.
+However, the platform has also been plagued by bots and fake accounts that are used to manipulate and spread false information.
 According to the research community, the usage of manipulation techniques implemented with the use of bot accounts is registered during diverse popular topic discussions.
-More specifically, studies show that bot accounts are involved in discussions around the 2016 and 2020 US Presidential elections (Golovchenko et al.
-2020; Badawy, Ferrara, and Lerman 2018; Howard, Kollanyi, and Woolley 2016; Shevtsov et al.
-2022, 2023b).
-Besides the US elections a high bot activity with spreading of misleading information is also detected during election periods (presidential/parliamentary/state) in countries like Germany, Sweden, France, Spain, etc.
-(Neudert, Kollanyi, and Howard 2017; Pastor-Galindo et al.
-2020; Bradshaw et al.
-2017; Fernquist, Kaati, and Schroeder 2018; Castillo et al.
-2019; Rossi et al.
-2020).
-Furthermore, election discourse is not the only point of interest for bot account creators.
-Spreading propaganda, advertisement, and fake news are identified during the vaccination debate (Broniatowski et al.
-2018), the advertisement of e-cigarette (Allem and Ferrara 2016; Allem et al.
-2017) and more recent examples of COVID-19 pandemic (Shahi, Dirkson, and Majchrzak 2021; Ferrara 2020; Yang et al.
-2021).
+More specifically, studies show that bot accounts are involved in diverse political discussions, for instance, around 2016 in countries like the US, Germany, Sweden, France, Spain, etc.
+(Golovchenko et al., 2020; Badawy et al., 2018; Howard et al., 2016; Shevtsov et al., 2022b, 2023; Neudert et al., 2017; Pastor-Galindo et al., 2020; Bradshaw et al., 2017; Fernquist et al., 2018; Castillo et al., 2019; Rossi et al., 2020).
+Furthermore, propagandizing and false news is identified during the vaccination debate (Broniatowski et al., 2018) and more recent examples of the COVID-19 pandemic (Shahi et al., 2021; Ferrara, 2020; Yang et al., 2021).
 This high activity of bot accounts raises concerns in the research community and online social media platforms about the integrity of the shared information.
-Twitter has been taking measures to detect and suspend these bots to maintain the integrity of the platform.
-Most of the published studies utilize advanced ML techniques to analyze the content and features (structural/network).
-Currently, the research community offers a variety of machine learning and neural network-based bot detection methods.
+Currently, the research community offers a variety of ML and NN-based bot detection methods.
 Each of these provided methods excels at addressing specific bot detection scenarios, yielding optimal solutions for particular use cases.
 Unfortunately, the performance of these provided methods significantly diminishes in more general bot detection scenarios, encompassing different periods, discussion topics, and languages, among others.
-Recent studies have demonstrated that existing methods still fall short of achieving flawless Twitter bot detection, regardless of the number of user characteristics they incorporate.
-In our research, we address this challenge by developing a semi-automatic machine learning pipeline (SAMLP) for constructing a generic bot detection model.
-Next, we compare this developed model with 34 currently available state-of-the-art approaches using nine publicly available datasets.
-This research yields two distinct evaluation scenarios: one specific to each dataset, where each model is trained and tested on each dataset separately, and a more generic approach where the models are trained and tested on the combined dataset comprising all nine unique datasets.
-As a result of our research, we introduce BotArtist, a finely tuned general bot detection model that achieves the highest F1-score performance on three out of the nine datasets, with an average F1-score performance of 83.19.
-Additionally, the presented approach achieves greater accuracy in the general evaluation, with an overall F1-score improvement of more than 9% and an average improvement of 5% compared to the best-performing existing bot detection approaches.
+Recent studies have shown that existing methods still do not achieve flawless Twitter bot detection, regardless of the number of user characteristics that they incorporate (Feng et al., 2022a).
+In addition, a large portion of existing ML bot detection methods ignore well-known optimization procedures (including feature selection and dynamic hyper-parameter fine-tuning) providing significant space for improvement.
+In our research, we address this challenge by developing a semi-automatic machine learning pipeline (SAMLP [^1] GitHub repository: TBD ) for constructing a generic bot detection model (BotArtist [^2] GitHub repository: TBD ).
+The presented pipeline resolves popular issues during the creation of the classification model, including recursive hyperparameter fine-tuning during the feature selection phase.
+This approach guarantees the reduction of non-relevant features which sequentially reduces the noise of initial data.
+Additionally, our pipeline considers the class imbalance during feature selection and model fine-tuning, making the model suitable for real-case applications.
+Next, we compare this developed model with 35 currently available state-of-the-art approaches using nine publicly available datasets.
+This research yields two distinct evaluation scenarios: one specific to each dataset, where each model is trained and tested separately on each of them, and a more generic approach, where the models are trained and tested on the combined dataset comprising all nine unique datasets.
+As a result of our research, we introduce BotArtist, a finely tuned general bot detection model that achieves the highest F1-score performance on three of the nine datasets, with an average F1-score performance of 83.19.
+Furthermore, the presented approach achieves greater precision in general evaluation, with an overall improvement in F1 score of more than 9% and an average improvement of 5% compared to the best-performing existing bot detection approaches.
 Moreover, BotArtist relies on only a limited set of profile features, allowing for the prediction of historical data independently of the Twitter API.
-Finally, we offer insights into the decision-making process of BotArtist through the use of the SHAP method.
-This approach enables us to explain the model’s decisions and highlight the distinctions between normal and bot accounts as identified by the developed model.
-Upon paper acceptance, we will make the SAMLP pipeline and the generated BotArtist model publicly accessible.
-Related Work
+Except from the source code of SAMLP and BotArtist, we make available the model outcome usable for research.
+For this purpose, we collect a large volume of users correlated to the already existing textual shared dataset and add an additional layer to textual information via user prediction.
+We then release 10.929.533 user profiles with BotArtist predictions  [^3] Zenodo repository: TBD , correlated to 127.275.386 tweets (Shevtsov et al., 2022a).
+2 Related Work
 Bot detection on Twitter poses a formidable challenging task, primarily due to the increasing sophistication of these bots.
@@
 Each category offers a distinct angle on bot detection, leveraging various extracted characteristics of users and their activities on the social media platform.
-Feature based
-Methods falling into this category perform feature engineering based on information extracted from Twitter user profiles and their activity patterns.
-Researchers utilize traditional machine learning or neural network classification algorithms for bot detection.
-These approaches employ different sets of user characteristics as feature sets, including user metadata (Kudugunta and Ferrara 2018), tweets (Miller et al.
-2014), description (Hayawi et al.
-2022), temporal patterns (Mazza et al.
-2019), and follow relationships (Feng et al.
-2021a).
-Moreover, some aim to enhance the scalability of feature-based methods (Yang et al.
-2020), discover unknown bot accounts using correlation-based techniques (Chavoshi, Hamooni, and Mueen 2016), and improve the trade-off between precision and recall in bot detection (Morstatter et al.
-2016).
-However, creators of bot accounts are increasingly aware of the features utilized by the research community, leading to novel bot implementations designed to evade detection based on known features (Cresci 2020).
-Consequently, existing feature-based methods face several challenges in the accurate detection of these new bot accounts (Feng et al.
-2021a).
-Text based
+2.1 Feature based
+Methods falling into this category apply feature engineering based on information extracted from Twitter user profiles and their activity patterns.
+Researchers use traditional machine learning or neural network classification algorithms for bot detection.
+These approaches employ different sets of user characteristics as feature sets, including user metadata (Kudugunta and Ferrara, 2018), tweets (Miller et al., 2014), user name (Beskow and Carley, 2019), description (Hayawi et al., 2022), temporal patterns (Mazza et al., 2019), and follow relationships (Feng et al., 2021a).
+Moreover, some of them aim to enhance the scalability of feature-based methods (Yang et al., 2020; Abreu et al., 2020), discover unknown bot accounts using correlation-based techniques (Chavoshi et al., 2016), and improve the trade-off between precision and recall in bot detection (Morstatter et al., 2016).
+However, the creators of bot accounts are increasingly aware of the features utilized by the research community, leading to novel bot implementations designed to evade detection based on known features (Cresci, 2020).
+Consequently, existing feature-based methods face several challenges in the accurate detection of these new bot accounts (Feng et al., 2021a).
+2.2 Text based
 Text-based methods rely on natural language processing (NLP) techniques for Twitter bot detection, extracting characteristics from posted content (tweets) and user profile descriptions.
-Approaches in this category include word embeddings (Wei and Nguyen 2019), recurrent neural networks (RNN) (Kudugunta and Ferrara 2018), attention mechanisms (Feng et al.
-2021a), and the adoption of pre-trained language models for encoding tweets (Dukić, Keča, and Stipić 2020).
-Researchers have also combined tweet representations with user profile characteristics (Cai, Li, and Zengi 2017), employed unsupervised machine learning models (Feng et al.
-2021a), and addressed multilingual context issues (Knauth 2019).
-Despite the extensive existing research studies and impressive performances in text-based approaches, novel bot accounts can still evade detection by sharing stolen content from genuine users (Cresci 2020).
-Additionally, recent work has demonstrated that relying solely on textual information is insufficient for robust and accurate bot detection (Feng et al.
-2021c).
-Graph based
+Approaches in this category include sequence fingerprint (Cresci et al., 2016), word embeddings (Wei and Nguyen, 2019), recurrent neural networks (RNN) (Kudugunta and Ferrara, 2018), attention mechanisms (Feng et al., 2021a), transformers (Guo et al., 2021a; Liu et al., 2019; Raffel et al., 2020) and the adoption of pre-trained language models for encoding tweets (Dukić et al., 2020).
+Researchers have also combined tweet representations with user profile characteristics (Cai et al., 2017; Efthimion et al., 2018; Kantepe and Ganiz, 2017; Miller et al., 2014; Varol et al., 2017; Kouvela et al., 2020; Lee et al., 2011; Echeverria et al., 2018), employed unsupervised machine learning models (Feng et al., 2021a), and addressed multilingual context issues (Knauth, 2019).
+Despite extensive existing research studies and impressive performance in text-based approaches, new bot accounts can still evade detection by sharing stolen content from genuine users (Cresci, 2020).
+Additionally, recent work has demonstrated that relying solely on textual information is insufficient for robust and accurate bot detection (Feng et al., 2021b).
+2.3 Graph based
 The graph-based category of bot detection methods combines geometric deep neural networks with graph analytics.
-Current implementations leverage techniques such as node centrality (Dehghan et al.
-2023), node representation learning (Pham et al.
-2022), graph neural networks (GNNs) (Ali Alhosseini et al.
-2019), and heterogeneous GNNs (Feng et al.
-2021c) to conduct graph-based bot detection.
-In recent research, authors have borrowed ideas from other categories to merge different approaches, such as combining graph and text-based methods (Guo et al.
-2021b; Yang, Ferrara, and Menczer 2022).
-Furthermore, novel GNN architectures have been proposed to exploit the heterogeneity in the Twitter network (Feng et al.
-2022a).
+Current implementations leverage techniques such as node centrality (Dehghan et al., 2023), node representation learning (Pham et al., 2022), graph neural networks (GNNs) (Ali Alhosseini et al., 2019; Moghaddam and Abbaspour, 2022), and heterogeneous GNNs (Feng et al., 2021b) to conduct graph-based bot detection.
+In recent research, the authors have borrowed ideas from other categories to merge different approaches, such as combining multiple methods (Guo et al., 2021b; Yang et al., 2022; Knauth, 2019; Beskow and Carley, 2020; Rodríguez-Ruiz et al., 2020; Magelinski et al., 2020; Yang et al., 2013; Kipf and Welling, 2016, 2016; Veličković et al., 2017; Lv et al., 2021).
+Furthermore, novel GNN architectures have been proposed to exploit heterogeneity in the Twitter network (Feng et al., 2022b).
 Generally, these approaches hold significant promise for Twitter bot detection.
@@
 Simplified feature-based approaches struggle to generalize and provide robust results, while more complex methods based on text or graph characteristics require intensive computational resources and large datasets for model development.
-These limitations, coupled with the recent announcement of the monetizing Twitter API (Twitter 2023), consist of many existing methods expensive to maintain or nearly impossible to operate daily.
-In this research, we address these challenges and present robust and accurate Twitter bot detection methods that rely on a lightweight set of features.
-Our methods require only a single user profile object request and can predict users using historical user objects (Twitter API v1.1 or v2) without additional Twitter API requests, thereby significantly reducing operational costs.
-Datasets
-For this research paper, we have collected nine well-known publicly available datasets (Feng et al.
-2021b).
-All the selected datasets already contain ground truth labels, primarily obtained through manual analysis or crowd-sourcing.
-To simplify, we have labeled the selected datasets as follows: C-15 (Cresci et al.
-2015), G-17 (Gilani et al.
-2017), C-17 (Cresci et al.
-2017b, a), M-18 (Yang et al.
-2020), C-S-18 (Cresci et al.
-2018, 2019), C-R-19 (Mazza et al.
-2019), B-F-19 (Yang et al.
-2019), TwiBot-20 (Feng et al.
-2021b), and TwiBot-22 (Feng et al.
-2022b).
-In table: 1, we present the information provided in each dataset, along with the volume of normal and bot accounts.
-Dataset C-15 G-17 C-17 M-18 C-S-18 C-R-19 B-F-19 TwiBot-20 TwiBot-22
-# Total User 5,301 2,484 14,368 50,538 13,276 693 518 229,580 1,000,000
-# Human 1,950 1,394 3,474 8,092 6,174 340 380 5,237 860,057
-# Bot 3,351 1,090 10,894 42,446 7,102 353 138 6,589 139,943
-# Total Tweet 2,827,757 0 6,637,615 0 0 0 0 33,488,192 88,217,457
-# Human Tweet 2,631,730 0 2,839,361 0 0 0 0 33,488,192 81,250,102
-# Bot Tweet 196,027 0 3,798,254 0 0 0 0 33,488,192 6,967,355
-# Graph Edges 7,086,134 0 6,637,615 0 0 0 0 33,488,192 170,185,937
-Table 1: Description of the selected datasets and the information contained in those datasets.
-Figure 1: Machine learning pipeline used for BotArtist model creation. Includes feature selection and model fine-tuning; each step is executed during separate K-Fold cross-validation.
-Methodology
+Furthermore, as shown in (Shevtsov et al., 2024) the complex embedding-based model tends to perfectly capture the training data patterns, but lacks generalizability and requires frequent retraining over multiple datasets to keep high performance.
+These limitations, coupled with the recent announcement of the Twitter API monetization (Twitter, 2023), make many existing methods expensive to maintain or nearly impossible to operate daily.
+In this study, we address these challenges and present robust and accurate Twitter bot detection methods, based on a lightweight set of features.
+Our methods require only a single user profile object (Twitter API v1.1 or v2) to provide user prediction without additional Twitter API requests, thereby significantly reducing operational costs.
+3 Datasets
+For this research paper, we collect nine well-known publicly available datasets (Feng et al., 2021c).
+All selected datasets already contain ground truth labels, primarily obtained through manual analysis or crowd-sourcing.
+For simplicity, we label the selected datasets as follows: C-15 (Cresci et al., 2015), G-17 (Gilani et al., 2017), C-17 (Cresci et al., 2017a, b), M-18 (Yang et al., 2020), C-S-18 (Cresci et al., 2018, 2019), C-R-19 (Mazza et al., 2019), B-F-19 (Yang et al., 2019), TwiBot-20 (Feng et al., 2021c), and TwiBot-22 (Feng et al., 2022a).
+In Table 1, we present the information provided in each dataset, along with the volume of normal and bot accounts.
+Furthermore, in collaboration with (Shevtsov et al., 2022a) collect 10.929.533 Twitter profiles correlated with 127.275.386 publicly available tweets related to the public discussion topic of the 2022 Russo-Ukrainian War.
+Our collection of user profiles is based on the monitoring of selected topics starting from February 23, 2022, till June 23, 2023.
+The shared datasets contain a set of extracted features in an anonymized form of a CSV file and contain only preprocessed numerical features to protect user information.
+The provided dataset also provides anonymized user IDs which are identically correlated with publicly available user tweet dataset (Shevtsov et al., 2022a).
+Table 1: Description of selected datasets and information contained in those datasets.
+Dataset C-15 G-17 C-17 M-18 C-S-18 C-R-19 B-F-19 TwiBot-20 TwiBot-22 # Total User 5,301 2,484 14,368 50,538 13,276 693 518 229,580 1,$0$ Human 1,950 1,394 3,474 8,092 6,174 340 380 5,237 860,057 # Bot 3,351 1,090 10,894 42,446 7,102 353 138 6,589 139,943 # Total Tweet 2,827,757 0 6,637,615 0 0 0 0 33,488,192 88,217,457 # Human Tweet 2,631,730 0 2,839,361 0 0 0 0 927,292 81,250,102 # Bot Tweet 196,027 0 3,798,254 0 0 0 0 1,072,496 6,967,355 # Graph Edges 7,086,134 0 6,637,615 0 0 0 0 33,488,192 170,185,937
+Figure 1: Machine learning pipeline used for BotArtist model creation including feature selection and model fine-tuning; each step is executed during separate K-Fold cross-validation.
+4 Methodology
 The proposed approach is based on the development of a semi-automatic machine learning pipeline (SAMLP).
-We have chosen this implementation due to its simplicity for further usage and its ability to avoid many trivial mistakes during data processing, feature selection, hyper-parameter fine-tuning, and model evaluation.
-The steps of the developed pipeline are presented in figure: 1, where the initial step involves data splitting.
+We choose this implementation due to its simplicity for further usage and its ability to avoid many trivial mistakes during the ML model development such as data processing, feature selection, hyper-parameter fine-tuning, and model evaluation.
+The steps of the developed pipeline are presented in Figure 1, where the initial step involves data splitting.
 During data splitting, it is crucial to maintain a class balance between the training/validation and testing data portions.
-For this purpose, we utilize a stratified data split with a ratio of 70% for training/validation and 30% for the testing data portions, respectively.
-This data separation allows us to keep exactly 70% of each class during the training, while preserving the natural class distribution differences, present in the dataset.
-The testing data portion is kept hidden and is only utilized during the final testing of the model.
-Feature Selection
-After the data splitting, our pipeline applies a feature selection procedure over the train/validation data portion.
-For this purpose, we use the K-Fold cross-validation approach, which allows us to utilize the same data portion for multiple trainings and evaluations.
-Although K-Fold cross-validation may lead to slightly over-optimistic performance estimations, in our case, it is not relevant since we use this procedure for alpha hyper-parameter fine-tuning of the Lasso model, and we are interested in identifying the best alpha parameter.
-Additionally, since the Lasso regression model does not perform well over imbalanced datasets, we also take into account the class imbalance of the train/validation dataset.
-When the class balance is not equal, we utilize under-sampling of the majority class, making the data perfectly balanced and improving the prediction performance of the Lasso model.
+For this purpose, we apply a stratified data split with a 70:30 ratio for training, validation, and testing portions respectively.
+Such an approach allows us to preserve the natural class distribution difference.
+The testing data portion is kept hidden and is only utilized during the final testing of the model to prevent information leakage between train and test portions
+4.1 Feature Selection
+After the data splitting, our pipeline applies a feature selection procedure over the train/validation data portion during the K-Fold cross validation.
+Although K-Fold cross-validation may lead to slightly over-optimistic performance estimations, in our case, it is not relevant since we use this procedure for $\alpha$ hyper-parameter fine-tuning of the Lasso model, and we are interested in identifying the best $\alpha$ parameter.
+Additionally, since the Lasso regression model does not perform well over imbalanced datasets, we also take into account the class imbalance of the dataset.
+In case of data imbalance,we utilize under-sampling of the majority class, making the data perfectly balanced and improving the prediction performance of the Lasso model.
 Due to the high-class imbalance and the under-sampling of the majority class, there is a very high probability of losing some important samples, which may lead to inaccurate feature selection.
-To address this, we utilize 10 executive repetitions to capture as many samples of the majority class as possible.
-During these repetition rounds, we store the best alpha parameters based on the square mean error metric from the entire stratified k-Fold cross-validation results, where $K=5$.
+To address this, we utilize 10 consecutive repetitions to capture as many samples of the majority class as possible.
+During these repetition rounds, we store the best $\alpha$ parameters based on the square mean error metric from the entire stratified k-Fold cross-validation results, where $K=5$.
 At the end of the procedure, we compute the most frequent best $\alpha$ parameter and train the Lasso model with the entire train/validation dataset, keeping features with non-zero coefficients.
 Additionally, we check if the selected $\alpha$ parameter falls within the defined limits of the searching area.
-If it falls on the minimum or maximum values, we create a new searching area based on the original site to ensure the selected $\alpha$ is genuinely the best.
-This approach allows us to utilize feature selection without additional knowledge of the original data and modifications, as the procedure automatically finds the best alpha values, manages class balance and imbalance via under-sampling and repeated executions, and processes both perfectly balanced and highly imbalanced datasets without information loss.
-Model Fine-Tuning
+If it falls on the minimum or maximum values, we create a new search area based on the original site to ensure the selected $\alpha$ is genuinely the best.
+This approach allows us to utilize feature selection without additional knowledge of the original data and modifications, as the procedure automatically finds the best $\alpha$ values, manages class balance and imbalance via under-sampling and repeated executions, and processes both perfectly balanced and highly imbalanced datasets without information loss.
+4.2 Model Fine-Tuning
 Having reduced the problem’s dimensionality through feature selection, we now focus on identifying the model and its configuration (hyper-parameter set) that can provide accurate predictions over unseen data samples.
-We selected three well-known machine learning classification models: SVM (Boser, Guyon, and Vapnik 1992), RandomForest (Breiman 2001), and state-of-the-art XGBoost (Extreme Gradient Boosting) (Chen et al.
-2015).
-For each model, we defined large hyper-parameter ranges to cover a variety of configurations.
-To reduce complexity and execution time, we developed a sampling method that randomly selects $C$ unique configurations per model (in our case, $C=50$).
+We select three well-known machine learning classification models: SVM (Boser et al., 1992), RandomForest (Breiman, 2001), and state-of-the-art XGBoost (Extreme Gradient Boosting) (Chen et al., 2015).
+For each model, we define large hyper-parameter ranges to cover a variety of configurations.
+To reduce complexity and execution time, we develop a sampling method that randomly selects $C$ unique configurations per model (in our case, $C=50$).
 This method allows us to estimate different possible configurations without sacrificing execution time.
@@
 This model configuration becomes the final model, which will be evaluated over the hold-out (testing) data portion.
-In the final stage of the model development, we provide model explainability using the SHAP game-theoretical approach (Shapley 1953).
+In the final stage of the model development, we provide model explainability using the SHAP game-theoretical approach (Shapley, 1953).
 This allows us to describe the reasons behind the model’s decisions and highlight the distinctions between classes.
@@
 The developed model is then trained over the entire dataset, including the train/validation and testing sets, and can be utilized for real-world applications with the identified decision threshold.
-Feature Type Feature Type calculation
-name_len count screen_name_sim real-valued Jaccard of (screen_name, user_name)
-screen_name_len count foll_friends real-valued follower / friends
-description_len count age real-valued created_at / collection date
-listed count listed_by_age real-valued listed / account age
-statuses count statuses_by_age real-valued statuses / account age
-followers count followers_by_age real-valued followers / account age
-following count following_by_age real-valued friends / account age
-name_upper_len count name_upper_pcnt real-valued percentage of upper case
-name_lower_len count name_lower_pcnt real-valued percentage of lower case
-name_digits_len count name_digits_pcnt real-valued percentage of digits
-name_special_len count name_special_pcnt real-valued percentage of other characters
-screen_name_upper_len count screen_name_upper_pcnt real-valued percentage of upper case
-screen_name_lower_len count screen_name_lower_pcnt real-valued percentage of lower case
-screen_name_digits_len count screen_name_digits_pcnt real-valued percentage of digits
-screen_name_special_len count screen_name_special_pcnt real-valued percentage of other characters
-description_upper_len count description_upper_pcnt real-valued percentage of upper case
-description_lower_len count description_lower_pcnt real-valued percentage of lower case
-description_digits_len count description_digits_pcnt real-valued percentage of digits
-description_special_len count description_special_pcnt real-valued percentage of other characters
-description_urls count name_entropy real-valued entropy of user name
-description_mentions count screen_name_entropy real-valued entropy of the screen name
-description_hashtags count has_location boolean
-total_urls count has_profile_image boolean
-protected boolean has_profile_url boolean
-verified boolean
 Table 2: List of the profile features extracted from the Twitter API user objects.
-Method Type C-15 G-17 C-17 M-18 C-S-18 C-R-19 B-F-19 TB-20 TB-22 Average
-SGBot F 77.9 72.1 94.6 99.5 82.3 82.7 49.6 84.9 36.6 75.57
-Kudugunta et al. F 75.3 49.8 91.7 94.5 50.9 49.2 49.6 47.3 51.7 62.22
-Hayawi et al. F 85.6 34.7 93.8 91.5 60.8 60.9 20.5 77.1 24.7 61.06
-BotHunter F 97.2 69.2 91.6 99.6 82.2 82.9 49.6 79.1 23.5 74.98
-NameBot F 83.4 44.8 85.7 91.6 61.1 67.5 38.5 65.1 0.5 59.80
-Abreu et al. F 76.4 66.7 95.0 97.9 76.9 83.5 53.8 77.1 53.4 75.63
-BotArtist F 98.3 76.1 97.0 99.7 80.6 88.3 68.4 82.2 58.2 83.19
-Cresci T 1.17 - 22.8 - - - - 13.7 - -
-Wei T 82.7 - 78.4 - - - - 57.3 53.6 -
-BGSRD T 90.8 35.7 86.3 90.5 58.2 41.1 13.0 70.0 21.1 56.30
-RoBERTa T 95.8 - 94.3 - - - - 73.1 20.5 -
-T5 T 89.3 - 92.3 - - - - 70.5 20.2 -
-Efthimion FT 94.1 5.2 91.8 95.9 68.2 71.7 0.0 67.2 27.5 57.95
-Kantepe FT 78.2 - 79.4 - - - - 62.2 58.7 -
-Miller FT 83.8 59.9 86.8 91.1 56.8 43.6 0.0 74.8 45.3 60.23
-Varol FT 94.7 - - - - - - 81.1 27.5 -
-Kouvela FT 98.2 66.6 99.1 98.2 80.4 81.1 28.1 86.5 30.0 74.24
-Santos FT 78.8 14.5 83.0 92.4 65.2 75.7 21.0 60.3 - -
-Lee FT 98.6 67.8 99.3 97.9 82.5 82.7 50.3 80.0 30.4 76.61
-LOBO FT 98.8 - 97.7 - - - - 80.8 38.6 -
-Moghaddam FG 73.9 - - - - - - 79.9 32.1 -
-Alhosseini FG 92.2 - - - - - - 72.0 38.1 -
-Knauth FTG 91.2 39.1 93.4 91.3 94.0 54.2 41.3 85.2 37.1 69.64
-FriendBot FTG 97.6 - 87.4 - - - - 80.0 - -
-SATAR FTG 95.0 - - - - - - 86.1 - -
-Botometer FTG 66.9 77.4 96.1 46.0 79.6 79.0 30.8 53.1 42.8 63.5
-Rodrifuez-Ruiz FTG 87.7 - 85.7 - - - - 63.1 56.6 -
-GraphHist FTG 84.5 - - - - - - 67.6 - -
-EvolveBot FTG 90.1 - - - - - - 69.7 14.1 -
-Dehghan FTG 88.3 - - - - - - 76.2 - -
-GCN FTG 97.2 - - - - - - 80.8 54.9 -
-GAT FTG 97.6 - - - - - - 85.2 55.8 -
-HGT FTG 96.9 - - - - - - 88.2 39.6 -
-SimpleHGN FTG 97.5 - - - - - - 88.2 45.4 -
-BotRGCN FTG 97.3 - - - - - - 87.3 57.5 -
-RGT FTG 97.8 - - - - - - 88.0 42.9 -
-Table 3: The performance of each selected bot detection model, as reported in the (Feng et al.
-2021b) paper, is compared with that of BotArtist.
+Feature Type Feature Type Calculation name_len count screen_name_sim real-valued Jaccard of (screen_name, user_name) screen_name_len count foll_friends real-valued follower / friends description_len count age real-valued created_at / collection date listed count listed_by_age real-valued listed / account age statuses count statuses_by_age real-valued statuses / account age followers count followers_by_age real-valued followers / account age following count following_by_age real-valued friends / account age name_upper_len count name_upper_pcnt real-valued percentage of upper case name_lower_len count name_lower_pcnt real-valued percentage of lower case name_digits_len count name_digits_pcnt real-valued percentage of digits name_special_len count name_special_pcnt real-valued percentage of other characters screen_name_upper_len count screen_name_upper_pcnt real-valued percentage of upper case screen_name_lower_len count screen_name_lower_pcnt real-valued percentage of lower case screen_name_digits_len count screen_name_digits_pcnt real-valued percentage of digits screen_name_special_len count screen_name_special_pcnt real-valued percentage of other characters description_upper_len count description_upper_pcnt real-valued percentage of upper case description_lower_len count description_lower_pcnt real-valued percentage of lower case description_digits_len count description_digits_pcnt real-valued percentage of digits description_special_len count description_special_pcnt real-valued percentage of other characters description_urls count name_entropy real-valued entropy of user name description_mentions count screen_name_entropy real-valued entropy of the screen name description_hashtags count has_location boolean total_urls count has_profile_image boolean protected boolean has_profile_url boolean verified boolean
+Table 3: The performance of each selected bot detection model, as reported in the (Feng et al., 2021c) paper, is compared with that of BotArtist.
 Performance is measured using the F1-score.
 In this benchmark, each model is trained and tested on each dataset separately.
-Method C-15 G-17 C-17 M-18 C-S-18 C-R-19 B-F-19 TB-20 TB-22 Total Average
-BotArtist 82.7 39.9 87.3 99.0 80.6 73.8 16.6 80.3 56.9 63.7 68.5
-Lee 82.3 0.0 83.6 97.7 78.2 67.7 20.0 8.5 42.4 52.9 53.3
-Abreu 84.4 0.3 80.1 88.4 67.1 40.8 11.7 15.6 29.0 40.4 46.3
-SGBot 75.0 3.6 79.8 99.2 76.7 68.9 0.0 15.2 43.3 53.8 51.5
-BotHunter 73.4 7.1 76.0 99.2 76.0 44.8 11.1 14.7 28.0 43.1 47.8
-Kouvela 95.5 20.4 94.7 98.1 78.4 71.4 21.0 28.5 36.0 52.0 60.5
-Botometer 66.9 77.4 96.1 46.0 79.6 79.0 30.8 53.1 42.8 45.3 63.5
-Table 4: The measurement of usage, like general bot detection approaches, involves training and testing models on all datasets. Performance is assessed using the F1-score.
-Feature Extraction
-As mentioned earlier, in this study we pursue the creation of a simple model that in comparison with other bot detection systems does not require a large volume of information or features for accurate prediction.
+Method Type C-15 G-17 C-17 M-18 C-S-18 C-R-19 B-F-19 TB-20 TB-22 Average SGBot F 77.9 72.1 94.6 99.5 82.3 82.7 49.6 84.9 36.6 75.57 Kudugunta et al.
+F 75.3 49.8 91.7 94.5 50.9 49.2 49.6 47.3 51.7 62.22 Hayawi et al.
+F 85.6 34.7 93.8 91.5 60.8 60.9 20.5 77.1 24.7 61.06 BotHunter F 97.2 69.2 91.6 99.6 82.2 82.9 49.6 79.1 23.5 74.98 NameBot F 83.4 44.8 85.7 91.6 61.1 67.5 38.5 65.1 0.5 59.80 Abreu et al.
+F 76.4 66.7 95.0 97.9 76.9 83.5 53.8 77.1 53.4 75.63 BotArtist F 98.3 76.1 97.0 99.7 80.6 88.3 68.4 82.2 58.2 83.19 Cresci T 1.17 - 22.8 - - - - 13.7 - - Wei T 82.7 - 78.4 - - - - 57.3 53.6 - BGSRD T 90.8 35.7 86.3 90.5 58.2 41.1 13.0 70.0 21.1 56.30 RoBERTa T 95.8 - 94.3 - - - - 73.1 20.5 - T5 T 89.3 - 92.3 - - - - 70.5 20.2 - Efthimion FT 94.1 5.2 91.8 95.9 68.2 71.7 0.0 67.2 27.5 57.95 Kantepe FT 78.2 - 79.4 - - - - 62.2 58.7 - Miller FT 83.8 59.9 86.8 91.1 56.8 43.6 0.0 74.8 45.3 60.23 Varol FT 94.7 - - - - - - 81.1 27.5 - Kouvela FT 98.2 66.6 99.1 98.2 80.4 81.1 28.1 86.5 30.0 74.24 Santos FT 78.8 14.5 83.0 92.4 65.2 75.7 21.0 60.3 - - Lee FT 98.6 67.8 99.3 97.9 82.5 82.7 50.3 80.0 30.4 76.61 LOBO FT 98.8 - 97.7 - - - - 80.8 38.6 - Moghaddam FG 73.9 - - - - - - 79.9 32.1 - Alhosseini FG 92.2 - - - - - - 72.0 38.1 - Knauth FTG 91.2 39.1 93.4 91.3 94.0 54.2 41.3 85.2 37.1 69.64 FriendBot FTG 97.6 - 87.4 - - - - 80.0 - - SATAR FTG 95.0 - - - - - - 86.1 - - Botometer FTG 66.9 77.4 96.1 46.0 79.6 79.0 30.8 53.1 42.8 63.5 Rodrifuez-Ruiz FTG 87.7 - 85.7 - - - - 63.1 56.6 - GraphHist FTG 84.5 - - - - - - 67.6 - - EvolveBot FTG 90.1 - - - - - - 69.7 14.1 - Dehghan FTG 88.3 - - - - - - 76.2 - - GCN FTG 97.2 - - - - - - 80.8 54.9 - GAT FTG 97.6 - - - - - - 85.2 55.8 - HGT FTG 96.9 - - - - - - 88.2 39.6 - SimpleHGN FTG 97.5 - - - - - - 88.2 45.4 - BotRGCN FTG 97.3 - - - - - - 87.3 57.5 - RGT FTG 97.8 - - - - - - 88.0 42.9 -
+Table 4: The measurement of performance in the case of general bot detection approaches, involves training and testing models on all datasets. Performance is assessed using the F1-score.
+Method C-15 G-17 C-17 M-18 C-S-18 C-R-19 B-F-19 TB-20 TB-22 Total Average BotArtist 82.7 39.9 87.3 99.0 80.6 73.8 16.6 80.3 56.9 63.7 68.5 Lee 82.3 0.0 83.6 97.7 78.2 67.7 20.0 8.5 42.4 52.9 53.3 Abreu 84.4 0.3 80.1 88.4 67.1 40.8 11.7 15.6 29.0 40.4 46.3 SGBot 75.0 3.6 79.8 99.2 76.7 68.9 0.0 15.2 43.3 53.8 51.5 BotHunter 73.4 7.1 76.0 99.2 76.0 44.8 11.1 14.7 28.0 43.1 47.8 Kouvela 95.5 20.4 94.7 98.1 78.4 71.4 21.0 28.5 36.0 52.0 60.5 Botometer 66.9 77.4 96.1 46.0 79.6 79.0 30.8 53.1 42.8 45.3 63.5
+4.3 Feature Extraction
+As mentioned earlier, in this study we pursue the creation of a simple model that in comparison with other bot detection systems- does not require a large volume of information or features for accurate prediction.
 In our case, the number of Twitter profiles triggers our attention.
 The majority of social media bot accounts are created for content promotion and increasing the attention of other registered users.
-To maximize this goal, bot accounts pursue the expansion of social audience.
+To maximize this goal, bot accounts aim for the expansion of social audience.
 Social expansion is very difficult to achieve without high activity, such as tweets, re-tweets, comments, etc.
-In comparison with profile-based features, the extraction and utilization of textual or graph features can result in instability, causing a significant performance drop across various periods or discussion topic datasets (Shevtsov et al.
-2023a).
+In comparison with profile-based features, the extraction and utilization of textual or graph features can result in instability, causing a significant performance drop across various periods or discussion topic datasets (Shevtsov et al., 2024).
 Consequently, we need to extract as much information as possible from the user profiles to identify extremely active accounts.
 The difference between very popular social media accounts (also known as celebrities), and bot accounts is that they try to post and share as much content as possible during the day.
-To properly separate normal and bot accounts, we develop 49 unique features described in table: 2.
 Due to the differences between Twitter API v1.1 and v2, we have devised a method for constructing features in a way that allows the extraction of selected information from both v1.1 and v2 without any additional manipulation.
@@
 The initial count feature categories encompass raw Twitter API object values, including followers, followings, status, and number of subscribed Twitter lists.
-In addition to these raw values, we have calculated statistics for textual fields such as user name, screen_name, and description.
-For these features, we have quantified the number of distinct characters, including uppercase letters, lowercase letters, digits, and special characters that do not fall into any of the previous categories.
-Furthermore, we have leveraged the profile description field, where users can provide additional information such as mentions of other users/organizations, hashtags, and additional URLs.
+In addition to these raw values, we calculate statistics for textual fields such as user name, screen_name, and description.
+For these features, we quantify the number of distinct characters, including uppercase letters, lowercase letters, digits, and special characters that do not fall into any of the previous categories.
+Furthermore, we leverage the profile description field, where users can provide additional information such as mentions of other users/organizations, hashtags, and additional URLs.
 In this context, we have counted the volume of the mentioned entities in the description field as raw values.
-In contrast to the count categories, where values are represented as integers, the real-valued category contains floating-point values.
-We have conducted various comparisons and measurements across the profile fields in this category.
+In contrast to the count categories, where values are represented as integers, the real-valued categories contain floating-point values.
+We conduct various comparisons and measurements across the profile fields in this category.
 Initially, we compute the age of the accounts, which allows us to distinguish highly active users with millions of posts spanning several years from users with similar activity but with an active period of only a few days.
@@
 Based on the account’s age, we can compute metrics similar to those in the count values category but adjusted for the account’s age perspective.
-For example, we compute metrics such as statuses_by_age, follower_by_age, and following_by_age using the formula:
-$feature\_by\_age=\frac{feature\ value}{profile\ age}$
 Given that many bot accounts are mass-created by automated scripts, they often have very similar or even trivial user names and screen names.
-Therefore, we have also applied the Jaccard similarity measure between user names and screen names.
-Jaccard similarity calculates the intersection of characters used in both fields in relation to the differences between them, using the following formula:
-$Jaccard\ similarity(A,B)=\frac{|A\cap B|}{|A\cup B|}$
-Additionally, we have computed the entropy of the user name and screen_name, enabling us to assess the randomness of these selected fields.
-In addition to the count values derived from the profile text fields (user name, screen_name, and description), we have computed the percentage of each specific character category, including lowercase, uppercase, digits, and special characters, relative to the length of the text field.
+Therefore, we apply the Jaccard similarity measure between user names and screen names to calculate the intersection of characters used in both fields to the differences between them.
+Additionally, we compute the entropy of the user name and screen_name, enabling us to assess the randomness of these selected fields.
+In addition to the count values derived from the profile text fields (user name, screen_name, and description), we compute the percentage of each specific character category, including lowercase, uppercase, digits, and special characters, relative to the length of the text field.
 These measurements offer a more precise means of identifying textual preferences more flexible to varying text lengths since they are computed as percentages of the field length.
@@
 For instance, we determine if the user account is protected, and verified, as well as extract information about the provided user location, profile URL, and whether the profile uses the default profile image.
-By employing this comprehensive feature engineering approach, we have maximized feature extraction while working with the limited profile information available through Twitter API v2.
-In total, we have extracted 49 unique profile features, representing the full set of features used before the feature selection process in the BotArtist model.
-Experimental Results
-According to the methodology of SAMLP described earlier, we’ve successfully developed a semi-automated machine learning pipeline capable of constructing classification models (both binary and multi-class) for both balanced and imbalanced datasets.
-Using this pipeline, we created and tested our custom implementation called BotArtist for Twitter bot detection, utilizing nine well-known datasets.
-To evaluate BotArtist’s performance, in comparison to 34 other existing research approaches, we employed the most comprehensive Twitter bot detection benchmark available (Feng et al.
-2022b), where various methods are compared using identical training/validation and testing data portions.
-This bench-marking allowed us to make a fair comparison between our approach and established bot detection methods.
+By employing this comprehensive feature engineering approach, we maximize the feature extraction while working with the limited profile information available through Twitter API v2.
+In total, we extract 49 unique profile features Table 2, representing the full set of features used before the feature selection process in the BotArtist model.
+5 Experimental Results
+According to the methodology of SAMLP described earlier, we successfully develop a semi-automated machine learning pipeline capable of constructing classification models (both binary and multiclass) for both balanced and imbalanced datasets.
+Using this pipeline, we create and test our custom implementation called BotArtist for Twitter bot detection, utilizing nine well-known datasets.
+To evaluate BotArtist’s performance, in comparison to 35 other existing research approaches, we employ the most comprehensive Twitter bot detection benchmark available (Feng et al., 2022a), where various methods are compared using identical training/validation and testing data portions.
+This benchmarking allows us to make a fair comparison between our approach and established bot detection methods.
 In our comparison, we not only evaluate BotArtist against bot detection approaches based on similar categories of features but also compare it to existing approaches developed based on different sets of characteristics.
-The selected methods are categorized into five different groups.
-Profile-Based features (F)
-This category includes approaches that rely on features extracted from user profile objects, including SGBot (Yang et al.
-2020), Kudugunta (Kudugunta and Ferrara 2018), Hayawi (Hayawi et al.
-2022), BotHunter (Beskow and Carley 2018), NameBot (Beskow and Carley 2019) and Abreu (Abreu, Ralha, and Gondim 2020) implementations.
-Text-Based features (T)
-In this category, we have approaches that focus on text characteristics, including Cresci (Cresci et al.
-2016), Wei (Wei and Nguyen 2019), BGSRD (Guo et al.
-2021a), RoBERTa (Liu et al.
-2019) and T5 (Raffel et al.
-2020).
-Profile and Text-Based features (FT)
-These approaches combine both profile and text-based features, and the selected implementations are Efthimion (Efthimion, Payne, and Proferes 2018), Kantepe (Kantepe and Ganiz 2017), Miller (Miller et al.
-2014), Varol (Varol et al.
-2017), Kouvela (Kouvela, Dimitriadis, and Vakali 2020), Lee (Lee, Eoff, and Caverlee 2011) and LOBO (Echeverria et al.
-2018).
-Profile and Graph-Based features (FG)
-Some approaches combine profile features with graph-based information. This category includes Moghaddam (Moghaddam and Abbaspour 2022) and Alhosseini (Ali Alhosseini et al. 2019) implementations.
-Profile, Text, and Graph-Based features (FTG)
-Approaches in this category aim to extract and combine as many features as possible, encompassing profile, text, and graph-based information.
-The selected implementations are Knauth (Knauth 2019), FriendBot (Beskow and Carley 2020), SATAR (Feng et al.
-2021a), Botometer (Yang, Ferrara, and Menczer 2022), Rodríguez-Ruiz (Rodríguez-Ruiz et al.
-2020), GraphHist (Magelinski, Beskow, and Carley 2020), EvolveBot (Yang, Harkreader, and Gu 2013), Dehghan (Dehghan et al.
-2023), GCN (Kipf and Welling 2016), GAT (Veličković et al.
-2017), HGT (Hu et al.
-2020), SimpleHGN (Lv et al.
-2021), BotRGCN (Feng et al.
-2021c) and RGT (Feng et al.
-2022a).
-Models comparison
-Our goal with BotArtist was to create a model with the highest possible level of generalizability.
-To achieve this, we designed two separate comparison scenarios.
-Initially, our focus was on assessing the fine-tuned BotArtist’s ability to capture general patterns within each dataset in our collection.
-As a result, we conducted training and testing for each of the selected models, including BotArtist, on each of the nine chosen datasets individually.
-This comparative approach allowed us to evaluate how well these selected models performed on each specific dataset, serving as a more rigorous bot detection implementation.
-Additionally, the designed experiments provided us with a broader understanding of the overall capabilities of the selected methods when applied in more general scenarios.
-While implementing these experiments, we took into account that the performance would likely be higher than that of a generic approach.
-Therefore, our primary focus was on discerning the differences between the models to select the most accurate bot detection methods for the broader, more generalized comparison.
+The selected methods are categorized into five different groups: feature-based (F), textual (T), graph (G), and any combination of them.
+5.1 Models comparison
+Our goal with BotArtist is to create a model with the highest possible level of generalizability.
+To achieve this, we design two separate comparison scenarios.
+Initially, our focus is on assessing the fine-tuned BotArtist’s ability to capture general patterns within each dataset in our collection.
+As a result, we conduct training and testing for each of the selected models, including BotArtist, on each of the nine chosen datasets individually.
+This comparative approach allows us to evaluate how well these selected models perform on each specific dataset, serving as a more rigorous bot detection implementation.
+Additionally, the designed experiments provide us with a broader understanding of the overall capabilities of the selected methods when applied in more general scenarios.
+While implementing these experiments, we take into account that the performance would likely be higher than that of a generic approach.
+Therefore, our primary focus is on discerning the differences between the models to select the most accurate bot detection methods for the broader, more generalized comparison.
+Based on the designed experiment we identified that (see Table 3) the BotArtist model emerged as the only bot detection method that achieves superior performance across three different datasets (M-18, C-R-19, B-F-19).
+Furthermore, the BotArtist model demonstrates the highest average F1-score, reaching 83.19.
+This represents a substantial improvement of 6.5% over the most accurate of the existing methods.
+As shown in Table 3, some datasets posed challenges for more complex methods that rely on text and graph features.
+These methods struggle to utilize datasets without text or graph information due to their limitations in accessing essential information required for the model.
 Previously, we had evaluated and compared models on each isolated dataset and calculated their average performance.
 While this approach provided us with an initial insight into model performance across diverse datasets, it lacked an assessment of model generalizability.
-To address this, we designed additional comparisons where we assessed the model’s ability to capture general patterns across different datasets.
-In these cases, we combined the training data from the nine datasets into a single dataset for model training.
-Combining data from multiple datasets, each with varying periods, discussion topics, and communities, demanded a high level of model generalization to avoid overfitting.
-To measure the performance of the trained methods, we evaluated them by testing on portions of each dataset and also on the union of all the dataset testing portions.
-These approaches allowed us to not only identify the general performance of each bot detection method but also gain insights into performance variations across each specific dataset.
-Parameter Value
-max_depth
-learning_rate
-subsample
-colsample_bytree
-min_child_weight
-gamma
-reg_lambda
-n_estimators
-eval_metric
-tree_method
-predictor
-objective
-scale_pos_weight
-selected features
-Table 5: The following are the selected parameters of the final general BotArtist (XGBoost) model, along with the number of features chosen during the feature selection procedure.
-In the initial model comparison experiment, we compared selected methods and assessed their performance based on the F1-score.
-This choice of metric was motivated by the highly imbalanced nature of the datasets, where proper comparison necessitates measurement metrics that account for class imbalances.
-In table: 3, we present the results of this comparison.
-Notably, the BotArtist model emerged as the only bot detection method that achieved superior performance across three different datasets (M-18, C-R-19, B-F-19).
-Furthermore, the BotArtist model demonstrated the highest average F1-score, reaching 83.19.
-This represents a substantial improvement of 6.5% over the most accurate of the existing methods, showcasing the model’s capacity to accurately capture patterns within each distinct dataset.
-As shown in table: 3, some datasets posed challenges for more complex methods that rely on text and graph features.
-These methods struggled to utilize datasets without text or graph information due to their limitations in accessing essential information required for the model.
-In the second model comparison scenario, our objective was to assess the generalization capabilities of various models.
-To achieve this, we selectively choose the most accurate models from table: 3, in addition to the Botometer model (Yang, Ferrara, and Menczer 2022).
-The Botometer model was included because it had already been trained on a diverse range of datasets and was expected to perform well in general-case scenarios.
-Additionally, the SAMLP pipeline manages to fine-tune BotArtists with the parameters presented in the table: 5 where the XGBoost classification model achieves higher performance during the K-Fold cross-validation.
-The results, as presented in table: 4, indicate that the most generalizable models are BotArtist, Botometer, and SGBot.
-BotArtist demonstrated superior performance over existing methods in terms of both F1-total and average scores, surpassing almost 10% existing methods in total.
+To address this, we design additional comparisons where we assess the model’s ability to capture general patterns across different datasets.
+In these cases, we combine the training data from the nine datasets into a single dataset for model training.
+Combining data from multiple datasets, each with varying periods, discussion topics, and communities, demands a high level of model generalization to avoid overfitting.
+To measure the performance of the trained methods, we evaluate them by testing on portions of each dataset and also on the union of all the dataset testing portions.
+These approaches allow us to not only identify the general performance of each bot detection method but also gain insights into performance variations across each specific dataset.
+To achieve this, we selectively choose the most accurate models from Table 3, in addition to the Botometer model (Yang et al., 2022).
+The Botometer model is included because it has already been trained on a diverse range of datasets and is expected to perform well in general-case scenarios.
+The results, as presented in Table 4, indicate that the most generalizable models are BotArtist, Botometer, and SGBot.
+BotArtist demonstrates superior performance over existing methods in terms of both total and average scores, surpassing almost 10% the existing methods in total.
 These results confirm that fine-tuned models which are based on the limited set of features are capable of capturing the general differences between normal and bot accounts.
-Furthermore, it’s noteworthy that the BotArtist model outperformed existing methods in both comparison scenarios.
-Due to the high volume of experiments, we do not provide any information according to the selected features and final model configuration for each evaluation scenario.
-These data will be shared and described in detail, in a GitHub repository after the paper acceptance.
+6 Conclusions and Future Work
+In this research paper, we introduce a semi-automatic machine learning pipeline (SAMLP) that addresses multiple challenges in creating machine learning models, including feature selection, hyperparameter fine-tuning, model evaluation, decision threshold optimization for binary classification, and provides model explainability through the SHAP game-theoretical approach.
+We apply this developed pipeline to create BotArtist, a versatile bot detection model based on user profile features.
+Our approach is trained and evaluated alongside current state-of-the-art solutions using nine different datasets.
+As demonstrated in our experiments, BotArtist surpasses existing, more complex methods in terms of generalization, achieving almost a 10% increase in the total F1 score in comparisons of multiple data sets and a 6.
+5% increase in comparisons of separated data sets.
+Additionally, we offer insights into the final model’s decision-making process and the patterns it captures, based on the SHAP game-theoretical approach (available at the repository).
+One of the limitations of the presented methodology is the relatively limited feature set, which might make it susceptible to evasion by future bot accounts that specifically aim to avoid detection based on these features.
+Further research conducted over a longer period is required to assess the model’s effectiveness in detecting bot accounts in the evolving landscape of social networks.
+For further research, we provide one of the largest labeled datasets containing user profiles correlated with another publicly available dataset of the 127M user posts.
+Such large labeled data will allow the development of state-of-the-art LLM models to detect bot-generated content in the post-Twitter API era.
+Appendix A Appendix
+A.1 Model benchmarks
+As mentioned in section 5 we compare BotArtist implementation with the existing 35 bot detection methods.
+We replicate the experiments of the model benchmarks described in (Feng et al., 2022a) where the authors provide the implementation of the 34 models.
+The Botometer (Yang et al., 2022) is an available model on-line that does not require additional training.
+Furthermore, the authors provide the original dataset split for training, validation, and testing portions, which allows the replication of the experiments with unbiased configurations.
+We manage to re-train each model (except the Botometer) over the training/validation set and evaluate their performance over the test set, as provided by (Feng et al., 2022a) (see Table 3).
+We differentiate from the original paper in the generic case scenario, as we combine the training/validation portions of the nine datasets into a single train portion where the selected top-performance models were trained over the combination of the datasets to evaluate the generalization of the methods.
+To reproduce the experiments, the repository  [^4] GitHub repository: TBD contains all the necessary scripts and information.
+The designed experiments were executed on a Dell server machine with two AMD EPYC 7543 CPUs (32 cores/64 threads), 512GB DDR4 RAM (16X32 GB) and nVidia A30 GPUs (24 GB VRAM), with a total execution time around 30 days.
+A.2 Created dataset
+Table 5: Set of hashtags used in our data collection query written in Russian, Ukrainian, and English provided with translation for non-English hashtags.
+#Ukraine, #Ukraina, #ukraina, #Украина(Ukraine), #Украине(Ukraine), #PrayForUkraine, #UkraineRussie, #StandWithUkraine, #StandWithUkraineNOW, #RussiaUkraineConflict, #RussiaUkraineCrisis, #RussiaInvadedUkraine, #WWIII, #worldwar3, #Война(War), #BlockPutinWallets, #UkraineRussiaWar, #Putin, #Russia, #Россия(Russia), #StopPutin, #StopRussianAggression, #StopRussia, #Ukraine_Russia, #Russian_Ukrainian, #SWIFT, #NATO, #FuckPutin, #solidarityWithUkraine, #PutinWarCriminal, #PutinHitler, #BoycottRussia, #with_russia, #FUCK_NATO, #ЯпротивВойны(I’m against war), #StopNazism #myfriendPutin #UnitedAgainstUkraine #StopWar #ВпередРоссия(Go Russia), #ЯМыРоссия(I/we Russia), #ВеликаяРоссия(Great Russia), #Путинмойпрезидент(Putin is my president), #россиявперед(Go Russia), #россиявперёд(Go Russia), #ПутинНашПрезидент(Putin is our president), #ЗаПутина(For Putin), #ПутинВведиВойска(Putin send the troops), #СЛАВАРОССИИ(Glory to Russia), #СЛАВАВДВ(Glory to Russian Airborne Forces)
+The provided 10M Twitter profiles dataset was retrieved through the Twitter API, while still available to the research community.
+The collection of the user profiles was initiated at the beginning of the 2022 Russo-Ukrainian War on 24 February 2022.
+We identify most of the popular hashtags based on the Ukrainian, Russian, and English hashtags during the initial days of invasion (see Table 5).
+Based on the collected hashtags we initiate the collection of the respective user profiles using them in their posts.
+Additionally, by using Twitter API we collect data from previous periods (23 February 2022) to capture some profiles and their activity 24 hours before the invasion.
+Furthermore, we collect on a daily basis the user profiles for the 16-month period until June 23, 2023.
+To share the collected user profiles and not violate user privacy and rights of the removal of their information from online social media, we provide the dataset in the processed form of CSV files where only numerical features are presented without any uuid information such as username, screen_name, description, URLs, user_id and profile_image_url.
+To hide the private user data while providing as much useful information for further usage in the case of ML/AI applications, we compute as many features as possible to maximize the information provided by the dataset (see Table 2).
+Additionally, based on the described BotArtist model we predict the collected user profiles and reveal which profile belongs to a human or bot account, based on BotArtist in combination with the model probability.
+Furthermore, we contact the original authors of the existing anonymized textual dataset over the same topic of Russo-Ukrainian War (Shevtsov et al., 2022a).
+We provide identical non-reversible anonymized uids to keep the connection between the user profile information and their tweets or user mentions over the 127.275.386 tweets.
+We believe that the combination of the provided labels and the large volume of textual data in the new era of Large Language Models will help the research community to develop state-of-the-art solutions for bot content detection.
+For this purpose the provided Zenodo dataset  [^5] Zenodo data repository: TBD and the metadata  [^6] GitHub repository: TBD are shared over the Attribution-NonCommercial-ShareAlike 4.0 International license (CC BY-NC-SA 4.0 DEED)  [^7] Dataset license: https://creativecommons.org/licenses/by-nc-sa/4.0/.
+Shared data will not be maintained due to updated Twitter API subscriptions and will be provided only in the current public version of Zenodo repository in form of CSV file (see data sample provided in Table 6 ).
+Table 6: Sample of the provided data with the entire feature set.
+age description_urls description_mentions description_hashtags screen_name_sim statuses statuses_by_age following following_by_age followers followers_by_age listed listed_by_age 786.0 0 0 0 0.2 1215 1.545801526717557 258 0.3282442748091603 80 0.1017811704834605 0 0.0 3410.0 0 0 0 0.4117647058823529 178614 52.37947214076247 95 0.0278592375366568 1642 0.481524926686217 41 0.0120234604105571 4266.0 2 0 0 0.8571428571428571 207601 48.66408813877168 8 0.0018752930145335 3911749 916.9594467885609 1895 0.4442100328176278 2083.0 0 0 0 0.6923076923076923 1271 0.6101776284205472 138 0.0662506000960153 30 0.0144023043686989 4 0.0019203072491598 5224.0 0 3 0 0.4166666666666667 27539 5.271630934150076 1394 0.2668453292496171 2006632 384.1179173047473 8880 1.6998468606431854 870.0 0 0 0 0.3076923076923077 2331 2.679310344827586 325 0.3735632183908046 152 0.1747126436781609 0 0.0 1239.0 1 2 0 1.0 33541 27.07102502017756 5 0.0040355125100887 2514015 2029.0677966101696 $2$ 1.7933817594834545 1328.0 0 0 0 0.0 10589 7.973644578313253 318 0.2394578313253012 193 0.1453313253012048 6 0.0045180722891566 1127.0 1 0 1 0.2666666666666666 1728 1.5332741792369122 44 0.0390417036379769 15908 14.115350488021296 213 0.188997338065661 2202.0 1 0 0 0.4 5059 2.297456857402361 110 0.0499545867393278 8959 4.068574023614896 6 0.0027247956403269 13.0 0 0 0 0.2105263157894736 746 57.38461538461539 0 0.0 49506 3808.153846153846 267 20.53846153846154 1288.0 0 0 0 1.0 237937 184.7336956521739 4758 3.694099378881988 7546 5.858695652173913 46 0.0357142857142857 4323.0 1 0 0 0.6153846153846154 39331 9.098080037011334 610 0.1411057136247976 26450 6.118436271108027 474 0.1096460791117279 3422.0 0 0 0 0.7142857142857143 182898 53.44769140853303 5271 1.540327293980129 5131 1.4994155464640562 5 0.0014611338398597 899.0 0 0 3 0.5714285714285714 8028 8.92992213570634 4864 5.410456062291435 3862 4.295884315906563 2 0.0022246941045606 925.0 0 0 0 0.8 401921 434.5091891891892 6203 6.705945945945946 9057 9.791351351351352 1 0.001081081081081 781.0 0 0 0 0.1333333333333333 17805 22.797695262483995 13313 17.0460947503201 17091 21.88348271446863 10 0.0128040973111395 4628.0 0 0 0 0.8181818181818182 7642 1.6512532411408816 1670 0.3608470181503889 87651 18.939282627484875 358 0.0773552290406223 4771.0 0 0 0 0.8181818181818182 27719 5.809893104171033 2558 0.5361559421504926 3283 0.6881156990148816 120 0.0251519597568643 profile_image description_upper_len description_lower_len description_digit_len description_spec_len description_upper_pcnt description_lower_pcnt description_digit_pcnt description_spec_pcnt description_len screen_name_upper_len screen_name_lower_len screen_name_digit_len 1 1 91 0 24 0.0086206896551724 0.7844827586206896 0.0 0.2068965517241379 116 0 12 0 1 0 0 0 153 0.0 0.0 0.0 1.0 153 0 13 1 1 10 34 0 68 0.0892857142857142 0.3035714285714285 0.0 0.6071428571428571 112 0 9 0 1 5 94 0 29 0.0390625 0.734375 0.0 0.2265625 128 3 12 0 1 22 93 0 38 0.1437908496732026 0.6078431372549019 0.0 0.2483660130718954 153 0 10 0 1 4 33 0 52 0.0449438202247191 0.3707865168539326 0.0 0.5842696629213483 89 2 2 3 1 12 72 1 36 0.0991735537190082 0.5950413223140496 0.0082644628099173 0.2975206611570248 121 1 9 0 1 0 35 0 7 0.0 0.8333333333333334 0.0 0.1666666666666666 42 0 10 0 1 26 75 2 30 0.1954887218045112 0.5639097744360902 0.0150375939849624 0.225563909774436 133 2 5 0 1 3 111 0 28 0.0211267605633802 0.7816901408450704 0.0 0.1971830985915492 142 0 11 0 1 43 18 6 13 0.5375 0.225 0.075 0.1625 80 2 0 5 1 3 85 0 23 0.027027027027027 0.7657657657657657 0.0 0.2072072072072072 111 1 7 3 1 14 102 4 23 0.0979020979020979 0.7132867132867133 0.0279720279720279 0.1608391608391608 143 0 14 0 1 3 47 0 17 0.044776119402985 0.7014925373134329 0.0 0.2537313432835821 67 1 5 2 1 4 69 0 28 0.0396039603960396 0.6831683168316832 0.0 0.2772277227722772 101 0 10 1 1 5 35 0 10 0.1 0.7 0.0 0.2 50 0 9 1 1 11 55 0 31 0.1134020618556701 0.5670103092783505 0.0 0.3195876288659793 97 2 9 0 1 14 88 2 22 0.1111111111111111 0.6984126984126984 0.0158730158730158 0.1746031746031746 126 0 14 0 1 15 105 0 39 0.0943396226415094 0.660377358490566 0.0 0.2452830188679245 159 0 10 0 screen_name_spec_len screen_name_upper_pcnt screen_name_lower_pcnt screen_name_digit_pcnt screen_name_spec_pcnt screen_name_len name_upper_len name_lower_len name_digit_len name_spec_len name_upper_pcnt name_lower_pcnt name_digit_pcnt 0 0.0 1.0 0.0 0.0 12 2 8 0 1 0.1818181818181818 0.7272727272727273 0.0 0 0.0 0.9285714285714286 0.0714285714285714 0.0 14 4 10 0 0 0.2857142857142857 0.7142857142857143 0.0 0 0.0 1.0 0.0 0.0 9 5 4 0 1 0.5 0.4 0.0 0 0.2 0.8 0.0 0.0 15 2 10 0 1 0.1538461538461538 0.7692307692307693 0.0 0 0.0 1.0 0.0 0.0 10 3 15 0 2 0.15 0.75 0.0 2 0.2222222222222222 0.2222222222222222 0.3333333333333333 0.2222222222222222 9 1 3 0 5 0.1111111111111111 0.3333333333333333 0.0 0 0.1 0.9 0.0 0.0 10 1 9 0 0 0.1 0.9 0.0 1 0.0 0.9090909090909092 0.0 0.0909090909090909 11 8 5 0 2 0.5333333333333333 0.3333333333333333 0.0 0 0.2857142857142857 0.7142857142857143 0.0 0.0 7 3 12 0 6 0.1428571428571428 0.5714285714285714 0.0 1 0.0 0.9166666666666666 0.0 0.0833333333333333 12 1 3 0 0 0.25 0.75 0.0 0 0.2857142857142857 0.0 0.7142857142857143 0.0 7 15 8 5 9 0.4054054054054054 0.2162162162162162 0.1351351351351351 0 0.0909090909090909 0.6363636363636364 0.2727272727272727 0.0 11 1 7 2 0 0.1 0.7 0.2 1 0.0 0.9333333333333332 0.0 0.0666666666666666 15 3 19 0 3 0.12 0.76 0.0 0 0.125 0.625 0.25 0.0 8 1 5 0 0 0.1666666666666666 0.8333333333333334 0.0 1 0.0 0.8333333333333334 0.0833333333333333 0.0833333333333333 12 0 14 0 1 0.0 0.9333333333333332 0.0 0 0.0 0.9 0.1 0.0 10 0 9 0 1 0.0 0.9 0.0 4 0.1333333333333333 0.6 0.0 0.2666666666666666 15 1 6 0 6 0.0769230769230769 0.4615384615384615 0.0 0 0.0 1.0 0.0 0.0 14 2 14 0 1 0.1176470588235294 0.8235294117647058 0.0 0 0.0 1.0 0.0 0.0 10 2 11 0 1 0.1428571428571428 0.7857142857142857 0.0 name_spec_pcnt name_len screen_name_entropy name_entropy has_location profile_url total_urls foll_friends verified protected uid label proba 0.0909090909090909 11 2.138333059508028 2.2718685126965625 1 1 1 0.3100775193798449 0 0 uidb46b… human 68 0.0 14 2.4410152780267027 2.4410152780267023 1 0 0 17.28421052631579 0 0 uida3ea… human 99 0.1 10 1.7351264569629226 2.302585092994045 1 1 3 488968.625 1 0 uidde3a… human 99 0.0769230769230769 13 2.5232109529528914 2.351673301904631 0 0 0 0.217391304347826 0 0 uid9dcc… bot 54 0.1 20 1.4708084763221112 2.4843668399738017 1 1 1 1439.477761836442 1 0 uid2991… human 99 0.5555555555555556 9 2.0431918705451206 2.19722457733622 1 0 0 0.4676923076923077 0 0 uida01b… human 95 0.0 10 2.0253262207700677 2.0253262207700677 1 1 2 502803.0 1 0 uid12f2… human 99 0.1333333333333333 15 1.9722469794234416 2.5232109529528914 1 1 1 0.6069182389937107 0 0 uidce17… human 93 0.2857142857142857 21 1.5498260458782016 2.532591160801121 1 1 2 361.54545454545456 0 0 uid8d37… human 91 0.0 4 2.253857589601352 1.3862943611198906 1 1 2 81.44545454545455 0 0 uid569e… human 93 0.2432432432432432 37 1.277034259466139 2.9147497445394843 1 0 0 0.0 0 0 uid5955… bot 92 0.0 10 2.0198149924929463 2.0253262207700677 0 0 0 1.585960487599832 0 0 uid5ef3… human 99 0.12 25 2.026229622995292 2.310717508000889 1 1 2 43.36065573770492 0 0 uidac7a… human 99 0.0 6 1.9061547465398496 1.5607104090414063 0 0 0 0.9734395750332006 0 0 uid059f… human 99 0.0666666666666666 15 2.2538575896013526 2.430791328878232 1 0 0 0.7939967105263158 0 0 uid55ca… human 94 0.1 10 2.1639556568820564 2.1639556568820564 0 0 0 1.460099951636305 0 0 uid9437… human 99 0.4615384615384615 13 2.2459520807289133 1.8446214763655029 1 1 1 1.2837827687223016 0 0 uid10f1… human 95 0.0588235294117647 17 2.106577333232116 2.3131536787336344 1 1 1 52.48562874251497 1 0 uid975d… human 99 0.0714285714285714 14 2.1639556568820564 2.4410152780267027 0 0 0 1.2834245504300237 0 0 uid4d4c… human 99
 Figure 2: BotArtist model decision explainability figure produced by SHAP methodology.
-Explainability
-Besides the creation of the high-performance model used for bot detection, we aim to shed light on the model decision itself.
-For this purpose we employ SHapley Additive exPlanations (SHAP) values (Lundberg and Lee 2017) which offer various benefits.
-One of the most significant advantages of SHAP values is that they are not tied to any particular type of machine learning model and are therefore model-agnostic.
-Furthermore, SHAP values display local accuracy, and consistency, which are characteristics that are not found concurrently in other methods.
-In this section we explain the BotArtist model trained over multiple datasets.
+A.3 Explainability
+This study aims to create a high performance model that can be used for bot detection while shedding light on the decision of the model itself.
+For this purpose, we employ SHapley Additive exPlanations (SHAP) values (Lundberg and Lee, 2017) which offer various benefits.
+One of the most significant advantages of SHAP values is that they are not tied to any particular type of machine learning model and, therefore, are model-agnostic.
+Furthermore, SHAP values display local accuracy and consistency and cannot be found concurrently in other methods.
+In this section, we explain the BotArtist model trained over multiple datasets.
 Figure: 2 presents the 20 most important features and their impact on the final decision.
@@
 Based on this order, we observe that the age of the account and the level of social activity are the most critical characteristics of our model, since the creation date of the account.
-The presented figure showcases the values of the features, ranging from low (indicated by the blue color), to high (indicated by the red color).
-Simultaneously, the x-axis provides information on the impact of each feature value on the model decision.
-Based on the presented results, it is crucial to highlight that the SHAP explanation technique provides insights that align with past related works.
-For instance, we find that bot accounts tend to have a significantly lower age of the account and lower number of followers combined with a higher number of followers by age of the account.
-Based on this information we can assume that bot accounts are not so popular in terms of total followers in comparison with normal users, but the growth of the profile followers is higher in comparison with normal users.
+The presented figure showcases the values of the features, ranging from low (indicated by the blue color) to high (indicated by the red color).
+The x-axis provides information on the impact of each feature value on the model decision.
+Based on the results presented, it is crucial to note that the SHAP explanation technique provides insights that align with past related work.
+For instance, we reveal that bot accounts tend to have a significantly lower age of the account and a lower number of followers combined with a higher number of followers by age of the account.
+Based on this information, we can assume that bot accounts are not so popular in terms of total followers in comparison with normal users, while the growth of the profile followers is higher in comparison with normal users.
 This assumption agrees with the original target of the bot account creators to achieve a higher audience during the short period.
-Conclusions and Future Work
-In this research paper, we introduce a semi-automatic machine learning pipeline (SAMLP) that addresses multiple challenges in creating machine learning models, including feature selection, hyperparameter fine-tuning, model evaluation, decision threshold optimization for binary classification, and provides model explainability through the SHAP game-theoretical approach.
-We apply this developed pipeline to create BotArtist, a versatile bot detection model based on user profile features.
-Our approach is trained and evaluated alongside current state-of-the-art solutions using nine different datasets.
-As demonstrated in our experiments, BotArtist surpasses existing, more complex methods in terms of generalization, achieving almost a 10% increase in total F1-score in multiple dataset comparisons and a 6.5% increase in separated dataset comparisons.
-Additionally, we offer insights into the final model’s decision-making process and the patterns it captures, based on the SHAP game-theoretical approach.
-Upon acceptance of this paper, we will make the developed SAMLP pipeline available for further research, along with the BotArtist model, which can label existing profile data.
-One of the limitations of the presented methodology is the relatively limited feature set, which might make it susceptible to evasion by future bot accounts that specifically aim to avoid detection based on these features.
-Further research conducted over a longer period is required to assess the model’s effectiveness in detecting bot accounts in the evolving landscape of social media.